### v1

# Emergence of Goal-Directed Behaviors via Active Inference with Self-Prior

Abstract
Infants often exhibit goal-directed behaviors, such as reaching for a sensory stimulus, even when no external reward criterion is provided. These intrinsically motivated behaviors facilitate spontaneous exploration and learning of the body and environment during early developmental stages. Although computational modeling can offer insight into the mechanisms underlying such behaviors, many existing studies on intrinsic motivation focus primarily on how exploration contributes to acquiring external rewards. In this paper, we propose a novel density model for an agent’s own multimodal sensory experiences, called the “self-prior,” and investigate whether it can autonomously induce goal-directed behavior. Integrated within an active inference framework based on the free energy principle, the self-prior generates behavioral references purely from an intrinsic process that minimizes mismatches between average past sensory experiences and current observations. This mechanism is also analogous to the acquisition and utilization of a body schema through continuous interaction with the environment. We examine this approach in a simulated environment and confirm that the agent spontaneously reaches toward a tactile stimulus. Our study implements intrinsically motivated behavior shaped by the agent’s own sensory experiences, demonstrating the spontaneous emergence of intentional behavior during early development. 00footnotetext: * Co-corresponding authors: Hoshinori Kanazawa, Yasuo Kuniyoshi 00footnotetext: Code is available at https://github.com/kim135797531/self-prior.
Research Highlights
– Suggests a computational model for early intentional behavior, integrating body-schema formation and goal-directed actions under the free energy principle.
– Introduces a self-prior as an internal density model that drives the emergence of goal-directed behaviors without external rewards.
– Demonstrates spontaneous reaching for a sticker by minimizing mismatches between observed multimodal sensory inputs and the empirically acquired self-prior.
1 Introduction
Infants spontaneously exhibit goal-directed behaviors, such as reaching for a sensory stimulus or actively exploring their surroundings, even in the absence of external rewards essential for survival. These behaviors are highly motivated by internal satisfaction and are referred to as intrinsically motivated behaviors (Czikszentmihalyi,, 1990; Ryan & Deci,, 2000). Such behaviors facilitate unsupervised learning without explicit external rewards or punishments, which is known to offer various advantages in early developmental stages (White,, 1959; Gopnik,, 2009; Zaadnoordijk et al.,, 2022; Kanazawa et al.,, 2023). For example, sensory experiences arising from an infant’s spontaneous movements, such as self-touch, not only promote learning about one’s own body (e.g., acquiring body representations) (Hoffmann et al.,, 2017), but also contribute to the formation of an early sense of self (Rochat,, 1998).
Traditional observational studies and neuroscience methodologies have been employed to investigate the mechanisms underlying intrinsically motivated behavioral development (Di Domenico & Ryan,, 2017). In particular, computational modeling has also been utilized to quantitatively interpret and predict behavioral development (Oudeyer & Kaplan,, 2009; Shultz,, 2013). Experiments using computer simulations allow researchers to freely manipulate and control specific variables to investigate various scenarios, making them advantageous for uncovering latent effects and interactions that may be difficult to observe through conventional behavioral studies alone.
Although computational modeling is a useful approach for studying infant behavioral development, many models proposed in modern robotics and machine learning tend to treat intrinsic motivation primarily as an exploratory mechanism for improving reward acquisition efficiency in environments with sparse external rewards (Aubret et al.,, 2023). Indeed, such approaches can enhance reinforcement learning agents’ ability to improve explicit task performance, such as maximizing video game scores. However, they do not sufficiently address how entirely new behaviors emerge when no explicit external reward objective is set, meaning that no performance criterion or information gain is provided.
To address this limitation, we propose a novel computational model of intrinsic motivation based on the expected free energy (Friston,, 2010; Friston et al.,, 2016) (Figure 1). We introduce an internal density model, termed the self-prior, which enables an agent to learn a statistical representation of incoming multimodal sensory signals. When a mismatch between the learned probabilistic model and actual observations is detected, goal-directed behavior emerges without any explicit reward criterion. Specifically, we implement a simulated agent that perceives a stimulus (such as a sticker) on its arm and spontaneously exhibits reaching behavior toward it. In the latter part of this paper, we discuss how the proposed self-prior mechanism may contribute to developmental science, neuroscience, and machine learning, positioning this study as a computational investigation into the early development of intentional behaviors in infants.
1.1 Free Energy Principle
Figure 1: Emergence of reaching behavior via the self-prior and active inference. (a) When a sticker is placed on the left arm of the simulated agent, it detects a mismatch with its prior experience of not having a sticker, and reaches toward the sticker with its right hand to minimize the discrepancy. (b) Development of the self-prior through experience: as sensory experiences are collected, the probability distribution over sensory patterns gradually develops. (c) The active inference process in which the agent plans future actions to minimize expected free energy by aligning sensory inputs with the learned self-prior. As a result, the agent performs a reaching action toward the sticker. A full-body infant illustration is used for clarity, the actual experiment was conducted in a pseudo-3D environment.
The free energy principle posits that an agent can learn to perceive and act by minimizing the "surprise" induced by sensory inputs (Friston,, 2010). In the free energy framework, surprise refers to the prediction error that arises when the agent encounters unexpected sensory observations, and reducing this surprise becomes the agent’s primary objective.
At each time step $t$, the agent receives information from the world through its sensory organs. We denote this sensory input by $o_{t}$, referring to it as the “observation.” Although $o_{t}$ is all the agent can directly observe, there may exist underlying factors in the world that generate these observations. Such factors are called hidden states, denoted by $s_{t}$ at time $t$. For example, given a properly learned model, an agent can infer from a tactile pattern on its arm (observation $o_{t}$) that there is a sticker attached to its arm (hidden state $s_{t}$).
Consider an agent that must satisfy certain biological criteria for survival in a stable manner. To do so, the agent must accurately understand its environment, which amounts to inferring $p(o_{t})$. According to the free energy principle, an agent maintains its own probabilistic model of these hidden states because doing so enables it to approximate $p(o_{t})$ and thus make inferences about the world. In practice, we often maximize the log of $p(o_{t})$ (i.e., $\log p(o_{t})$) for computational convenience, which is referred to as maximizing the model evidence. Minimizing the negative term, $-\log p(o_{t})$—known as surprise—is therefore equivalent to maximizing this model evidence.
In reality, calculating $p(o_{t})$ requires marginalization over hidden states (i.e., $p(o_{t},s_{t})$), which is often intractable. The variational free energy principle tackles this by introducing a variational approximation of Bayesian inference, wherein the agent indirectly reduces surprise by minimizing the variational free energy $\mathcal{F}$, an upper bound on $-\log p(o_{t})$. Formally, $\mathcal{F}$ is defined as follows (see (Parr et al.,, 2022) for details):
$-\log p(o_{t})\leq\mathcal{F}$ $=\mathbb{E}_{q(s_{t})}[\log q(s_{t})-\log p(o_{t},s_{t})]$ (1)
$=\underbrace{D_{\mathrm{KL}}[q(s_{t})\|p(s_{t})]}_{\text{complexity}}-\underbrace{\mathbb{E}_{q(s_{t})}[\log p(o_{t}\mid s_{t})]}_{\text{accuracy}}$
where $D_{\mathrm{KL}}(\cdot\|\cdot)$ denotes the Kullback–Leibler divergence, and $q(\cdot)$ represents the variationally approximated distribution. That is, $q(s_{t})$ is the variational distribution that approximates the posterior distribution over hidden states. Therefore, to reduce surprise at time $t$, the agent must keep the posterior and prior distributions close (reducing complexity) while maximizing the likelihood of observations (improving accuracy).
1.2 Active Inference and Expected Free Energy
While the free energy principle describes how an agent updates its model to reduce surprise based on current sensory observations, the agent can also directly change its observations through actions $a_{t}$ to make them better align its internal model. This is the essence of active inference, and more concretely, the approximate posterior $q(s_{t})$ is extended to include actions as $q(s_{t},a_{t})=q(a_{t}\mid s_{t})\,q(s_{t})$.
To obtain optimal actions $a_{t}$ for both present and future steps, a method has been proposed that computes expected free energy by considering the expected free energy at future time points (Friston et al.,, 2016). That is, the agent predicts multiple possible future states and selects the trajectory that is expected to minimize free energy most effectively.
However, since future observations cannot be obtained before the corresponding time actually arrives, the agent cannot directly compute the free energy of future states. Therefore, the expected free energy assumes that the agent possesses a preferred prior $\tilde{p}(o_{t})$ over desired observations, which acts as a setpoint. As a result, $p(o_{t},s_{t})$ is extended to include actions and is biased toward preferred observations, expressed as $\tilde{p}(o_{t},s_{t},a_{t})=p(a_{t}\mid s_{t})\,p(s_{t}\mid o_{t},a_{t})\,\tilde{p}(o_{t}\mid a_{t})$. Assuming $q(s_{t}\mid o_{t},a_{t})\approx p(s_{t}\mid o_{t},a_{t})$ and a uniform prior $p(a_{t})$ over actions, the expected free energy $\mathcal{G}$ is defined as follows (see (Millidge et al.,, 2020; Mazzaglia et al.,, 2021) for detailed approximations and derivation):
$\mathcal{G}$ $=\mathbb{E}_{q(o_{t},s_{t},a_{t})}[\log q(s_{t},a_{t})-\log\tilde{p}(o_{t},s_{t},a_{t})]$ (2)
$\approx-\underbrace{\mathbb{E}_{q(o_{t})}[\log\tilde{p}(o_{t})]}_{\text{extrinsic value}}-\underbrace{\mathbb{E}_{q(o_{t})}[D_{\mathrm{KL}}[q(s_{t}\mid o_{t})\|q(s_{t})]]}_{\text{intrinsic value}}-\underbrace{\mathbb{E}_{q(s_{t})}[\mathcal{H}(q(a_{t}\mid s_{t}))]}_{\text{action entropy}}$
where $\mathcal{H}$ denotes the information entropy. That is, to reduce surprise at a future time $t$, the agent must (i) increase the probability of preferred observations, (ii) obtain high information gain from those observations, and (iii) maintain diverse actions. As shown above, expected free energy integrates both pragmatic (extrinsic) and epistemic (intrinsic) value into a single expression, thereby naturally resolving the exploration–exploitation dilemma. The extrinsic motivation term is similar to a drive for achieving goals or obtaining rewards necessary for survival, whereas the intrinsic motivation term corresponds to an internal desire to explore new information or reduce uncertainty. Since both are measured in the same information unit (nats), they are automatically balanced within a unified system.
2 Self-Prior: Acquired Preference to Induce Goal-Directed Behavior
Under the expected free energy, agents do not require scalar rewards to be explicitly defined by the environment. This is because the agent is assumed to possess a distribution over preferred observations, called the preferred prior, which serves as the setpoint for action planning. However, since the preferred prior is typically given in a fixed form, the removal of external reward signals does not imply that the agent has autonomously generated its goal-directed behavior.
By contrast, in this study, we allow the agent to autonomously determine the setpoint for its action planning by additionally learning an internal density model, which we refer to as the self-prior. Through this, the agent spontaneously generates goal-directed behavior to minimize discrepancies between its internal model and the environment, including its own body.
Technically, we derive this by making a simple modification to the standard expected free energy formulation—separating the self-prior term from the original preferred prior term. This enables the extraction of an intrinsic motivation term that maintains preferences formed through experience history, as distinct from the extrinsic motivation term that maintains specific experiences. According to the classification proposed by (Oudeyer & Kaplan,, 2009), this corresponds to internal homeostatic intrinsic motivation, and in particular, is most closely related to what is termed familiarity-driven intrinsic motivation. We further examine in the Discussion section why intrinsic motivation can be extracted from the extrinsic motivation term in the free energy formulation.
To implement the self-prior in the active inference framework, we redefine each preferred observation $o$ in terms of two components: (i) an extrinsic observation $o^{E}$ that must be maintained to satisfy survival requirements, and (ii) an intrinsic observation $o^{I}$ related to the self-prior. Here, $o^{I}$ refers to bodily states or sensations—such as a tactile pattern on the arm—that can vary without threatening the agent’s essential functions. Consequently, we decompose $\tilde{p}(o)$ into $\tilde{p}(o^{E},o^{I})$, and in this study, we assume independence so that $\tilde{p}(o^{E},o^{I})=\tilde{p}(o^{E})\,\tilde{p}(o^{I})$. Thus, the expected free energy incorporating the self-prior becomes:
$\mathcal{G}$ $\approx-\underbrace{\mathbb{E}_{q(o_{t})}[\log\tilde{p}(o_{t})]}_{\text{extrinsic (homeostatic)}}-\underbrace{\mathbb{E}_{q(o_{t})}[D_{\mathrm{KL}}[q(s_{t}|o_{t})\|q(s_{t})]]}_{\text{intrinsic (heterostatic)}}-\mathcal{H^{\prime}}$ (3)
$=-\underbrace{\mathbb{E}_{q(o^{E}_{t},o^{I}_{t})}[\log\tilde{p}(o^{E}_{t},o^{I}_{t})]}_{\text{homeostatic}}-\underbrace{\mathbb{E}_{q(o_{t})}[D_{\mathrm{KL}}[q(s_{t}|o_{t})\|q(s_{t})]]}_{\text{heterostatic}}-\mathcal{H^{\prime}}$
$=-\underbrace{\mathbb{E}_{q(o^{E}_{t})}[\log\tilde{p}(o^{E}_{t})]}_{\text{extrinsic (homeo-, fixed)}}-\underbrace{\mathbb{E}_{q(o^{I}_{t})}[\log\tilde{p}(o^{I}_{t})]}_{\text{intrinsic (homeo-, familiarity)}}-\underbrace{\mathbb{E}_{q(o_{t})}[D_{\mathrm{KL}}[q(s_{t}|o_{t})\|q(s_{t})]]}_{\text{intrinsic (hetero-, novelty)}}-\mathcal{H^{\prime}}$
Here, $\mathcal{H}^{\prime}$ denotes the entropy of the action distribution.
We designed the system such that the agent can learn the self-prior autonomously, by automatically acquiring a mechanism that estimates familiarity based on the frequency of past observations. Details of the implementation are provided in the following section.
3 Computational Model of Self-Prior
In this section, we introduce the concrete implementation of our model for computational simulations based on the active inference framework incorporating the self-prior. We present separate models for environments with discrete and continuous variables. In the discrete environment, the model is built using small-sized matrices to allow direct inspection of its behavior. For the continuous environment, which requires a high-dimensional variable space, we use deep neural networks with techniques developed in deep reinforcement learning.
3.1 Discrete Model
The discrete model uses categorical distributions as the generative model, handling discrete random variables:
Prior: $P(s_{t}\mid s_{t-1},a_{t-1})=\text{Cat}(\mathbf{B}_{a_{t-1}})$ (4)
Posterior: $Q(s_{t})=\text{Cat}(\phi_{t})$
Likelihood: $P(o_{t}\mid s_{t})=\text{Cat}(\mathbf{A})$
Here, we use uppercase letters (e.g., $P,\,Q$) to indicate discrete distributions, and denote categorical distributions by $\text{Cat}(\cdot)$. The observation $o_{t}$ and hidden state $s_{t}$ are one-hot column vectors with $N_{o}$ and $N_{s}$ rows, respectively, corresponding to the number of possible observations and hidden states. Boldface symbols such as $\mathbf{A}$ and $\mathbf{B}$ indicate the matrix parameters of the categorical distributions. $\mathbf{B}$ contains separate matrices for each available action $a$ (e.g., if actions are {LEFT, STOP, RIGHT}, then separate matrices $\mathbf{B}_{\text{LEFT}},\mathbf{B}_{\text{STOP}},\mathbf{B}_{\text{RIGHT}}$ exist). $\mathbf{A}$ is an $N_{o}\times N_{s}$ matrix, and each $\mathbf{B}_{a_{t-1}}$ is an $N_{s}\times N_{s}$ square matrix. Based on these definitions, the variational free energy $\mathcal{F}$ under a given action $a_{t-1}$ can be rewritten as:
$\mathcal{F}$ $=\mathbb{E}_{Q(s_{t})}[\log Q(s_{t})-\log P(o_{t},s_{t})]$ (5)
$=\mathbb{E}_{Q(s_{t})}[\log{Q(s_{t})}-\log{P(s_{t})}]-\mathbb{E}_{Q(s_{t})}[\log P(o_{t}\mid s_{t})]$
$=\phi_{t}\cdot(\log{\phi_{t}}-\log{(\mathbf{B}_{a_{t-1}}\phi_{t-1})})-\phi_{t}\cdot(\log{(\mathbf{A}\cdot o_{t})})$
Accordingly, the parameter $\phi$ that minimizes free energy can be derived by taking the derivative of $\mathcal{F}$ with respect to $\phi$:
$\frac{\partial\mathcal{F}}{\partial\phi_{t}}$ $=\log{\phi_{t}}-\log{(\mathbf{B}_{a_{t-1}}\phi_{t-1})}+1-\log{(\mathbf{A}\cdot o_{t})}=0$ (6)
$\therefore\log{\phi_{t}}$ $\approx\log{(\mathbf{A}\cdot o_{t})}+\log{(\mathbf{B}_{a_{t-1}}\phi_{t-1})}$
$\phi_{t}$ $\approx\sigma(\log{(\mathbf{A}\cdot o_{t})}+\log{(\mathbf{B}_{a_{t-1}}\phi_{t-1})})$
Here, $\sigma$ denotes the softmax function, which normalizes the elements of $\phi$ so that they sum to 1. In other words, the new posterior parameter $\phi$ can be computed from the previous action $a_{t-1}$ and the current observation $o_{t}$.
We assume that the likelihood distribution $\text{Cat}(\mathbf{A})$ and the prior distribution $\text{Cat}(\mathbf{B})$ have already been learned accurately. That is, the parameter $\mathbf{A}$ perfectly specifies the observation $o_{t}$ for each possible hidden state $s_{t}$, and $\mathbf{B}$ perfectly specifies the outcome of each possible action $a_{t}$ for each state. This simplifying assumption allows us to isolate the effect of changes in the self-prior on behavior, without introducing side effects from other parts of the generative model (in the continuous model discussed later, the learning of both likelihood and prior is included).
The distribution $\tilde{P}(o_{t})$, which serves as the setpoint for future observations, is given in the form of a categorical distribution $\text{Cat}(\mathbf{C})$, where $\mathbf{C}$ is a column vector with a row for each possible $o_{t}$. In this study, since we aim to examine behavior driven by the self-prior, all $o_{t}$ contain only information related to $o^{I}_{t}$. Therefore, the entire $\text{Cat}(\mathbf{C})$ is defined as the self-prior:
Self-prior: $\tilde{P}(o^{I}_{t})=\text{Cat}(\mathbf{C})$ (7)
The agent autonomously learns the self-prior by recording the frequency of previously experienced observations and storing the corresponding observation probabilities in the parameter $\mathbf{C}$. The initial value of $\mathbf{C}$ is set so that all observations have equal probability.
To determine the optimal action $a_{t}$, we compute the expected free energy associated with each possible action, as described above. This is given by:
$\mathcal{G}$ $=\mathbb{E}_{Q(s_{t+1},o_{t+1}\mid a_{t})}[\log Q(s_{t+1}\mid a_{t})-\log\tilde{P}(o_{t+1},s_{t+1}\mid a_{t})]$ (8)
$=\mathbb{E}_{Q(s_{t+1},o_{t+1}\mid a_{t})}[\log Q(s_{t+1}\mid a_{t})-\log P(s_{t+1}\mid o_{t+1},a_{t})-\log\tilde{P}(o_{t+1})]$
$\approx\mathbb{E}_{Q(s_{t+1},o_{t+1}\mid a_{t})}[\log Q(s_{t+1}\mid a_{t})-\log Q(s_{t+1}\mid o_{t+1},a_{t})-\log\tilde{P}(o_{t+1})]$
$=\mathbb{E}_{Q(s_{t+1},o_{t+1}\mid a_{t})}[\log Q(o_{t+1}\mid a_{t})-\log Q(o_{t+1}\mid s_{t+1},a_{t})-\log\tilde{P}(o_{t+1})]$
For practical computation, the equation can be reformulated as follows:
$\mathcal{G}$ $\approx\mathbb{E}_{Q(s_{t+1}\mid a_{t})Q(o_{t+1}\mid s_{t+1},a_{t})}[-\log Q(o_{t+1}\mid s_{t+1},a_{t})]$ (9)
$\phantom{=======}+\mathbb{E}_{Q(o_{t+1}\mid a_{t})}[\log Q(o_{t+1}\mid a_{t})-\log\tilde{P}(o_{t+1})]$
$=\mathbb{E}_{Q(s_{t+1}\mid a_{t})P(o_{t+1}\mid s_{t+1})}[-\log P(o_{t+1}\mid s_{t+1})]$
$\phantom{=======}+\mathbb{E}_{Q(o_{t+1}\mid a_{t})}[\log Q(o_{t+1}\mid a_{t})-\log\tilde{P}(o_{t+1})]$
$=\mathbb{E}_{Q(s_{t+1}\mid a_{t})}[\mathcal{H}[P(o_{t+1}\mid s_{t+1})]]+D_{\mathrm{KL}}[Q(o_{t+1}\mid a_{t})\|\tilde{P}(o_{t+1})]$
$=(\mathbf{B}_{a_{t}}\phi_{t})\cdot\mathcal{H}[\mathbf{A}]+D_{\mathrm{KL}}[\mathbf{A}\mathbf{B}_{a_{t}}\phi_{t}\|\mathbf{C}]$
Here, we use $Q(o_{t}\mid s_{t},a_{t-1})=P(o_{t}\mid s_{t})$ because the mapping from state to observation is assumed to be independent of action. Furthermore, since $\mathbf{A}$ and $\mathbf{B}$ are assumed to be known exactly in our setup, the action entropy term for parameter information gain is omitted in the discrete experiment (see (Parr et al.,, 2022) for a detailed explanation of above formulation).
In the discrete environment, we compute expected free energy starting from the current time $t$ to a future time $t+N$. The total expected free energy for all steps is then used to evaluate a policy $\pi=\{a_{t},a_{t+1},...,a_{t+N}\}$. The policy is computed as a categorical distribution over actions, from which actions are sampled as follows:
Policy: $P(\pi)=\sigma(-\mathcal{G})$ (10)
3.2 Continuous Model
Prior research on extending the free energy principle from discrete to continuous domains often exploits the expressive power of deep neural networks. Indeed, several techniques originally proposed in deep reinforcement learning have been shown to closely approximate free energy optimization. For example, the Recurrent State-Space Model (RSSM) in PlaNet (Hafner et al.,, 2019) can be interpreted as optimizing variational free energy (Çatal et al.,, 2020), and policy gradient methods can be viewed as minimizing expected free energy (Millidge,, 2020). Following these precedents, we implement free energy minimization in a continuous environment using deep neural networks (Figure 2).
Whereas the discrete model fixed the parameters of the likelihood and prior distributions in variational free energy, the continuous model trains both distributions using deep neural networks, similar to the RSSM, with respect to the parameter $\phi$:
Prior: $p_{\phi}(h_{t}\mid s_{t-1},a_{t-1})$ (11)
Posterior: $q_{\phi}(s_{t}\mid h_{t},o_{t})$
Likelihood: $p_{\phi}(o_{t}\mid s_{t})$
The prior encodes the previous hidden state $s_{t-1}$ and action $a_{t-1}$ into a single vector via a linear layer, which is then fed into a GRU cell (Chung et al.,, 2014) to integrate inputs over time. The GRU’s hidden state is passed through a two-layer MLP that outputs the parameters of a multivariate Gaussian distribution, from which we sample a stochastic state. We concatenate this stochastic state with the GRU hidden state to produce the prior feature $h_{t}$. We use $h_{t}$ to distinguish the prior’s representation from the posterior variable $s_{t}$ (That is, the KL term in the variational free energy $\mathcal{F}$, originally $D_{\mathrm{KL}}[q(s_{t})\|p(s_{t})]$, is expressed as $D_{\mathrm{KL}}[q_{\phi}(s_{t}\mid h_{t},o_{t})\|p_{\phi}(h_{t}\mid s_{t-1},a_{t-1})]$).
(a) Original active inference
(b) Active inference with self-prior
Figure 2: Graphical model of active inference using deep neural networks that minimize variational free energy $\mathcal{F}$ and expected free energy $\mathcal{G}$. The self-prior $\tilde{p}(o^{I}_{t})$ is trained to maximize the log-likelihood of observations $o^{I}_{t}$. In the expected free energy calculation (highlighted in blue), the learned self-prior serves as the behavioral setpoint alongside the fixed preferred prior. Although the preferred prior and self-prior can theoretically be applied simultaneously, we use only the self-prior in this study for clarity of exposition; thus, the preferred prior is shown faded in the figure.
The posterior receives as input a concatenated vector of the prior feature $h_{t}$ and an embedded observation $o_{t}$. We use embedded observations because our experiment involves multimodal sensory data, and embedding allows us to integrate sensory inputs of different dimensionalities into a single representation. Since the embedding method may vary depending on the experiment, further details are provided in the Experiments section. The concatenated $h_{t}$ and $o_{t}$ are passed into a two-layer MLP in the posterior network, which outputs the parameters of a multivariate Gaussian distribution, and sample the posterior feature $s_{t}$ from this distribution.
The likelihood network transforms $s_{t}$ into $o_{t}$ through a linear layer, and the resulting embedded $o_{t}$ is then decoded into individual modality-specific observations. The parameters $\phi$ of the Prior, Posterior, and Likelihood networks are optimized via gradient descent to minimize the following free energy:
$\mathcal{F}$ $=\mathbb{E}_{q(s_{t})}[\log q(s_{t})-\log p(o_{t},s_{t})]$ (12)
$=\mathbb{E}_{q(s_{t})}[\log{q(s_{t})}-\log{p(s_{t})}]-\mathbb{E}_{q(s_{t})}[\log p(o_{t}\mid s_{t})]$
$=D_{\mathrm{KL}}[q(s_{t})\|p(s_{t})]-\mathbb{E}_{q(s_{t})}[\log p(o_{t}\mid s_{t})]$
$\therefore\operatorname*{\arg\min}_{\phi}\mathcal{F}$ $=\operatorname*{\arg\min}_{\phi}\big{[}D_{\mathrm{KL}}[q_{\phi}(s_{t}\mid h_{t},o_{t})\|p_{\phi}(h_{t}\mid s_{t-1},a_{t-1})]-\mathbb{E}_{q_{\phi}(s_{t}\mid h_{t},o_{t})}[\log p_{\phi}(o_{t}\mid s_{t})]\big{]}$
In the continuous model, as in the discrete case, we do not consider extrinsic preferences in order to focus on behavior driven by the self-prior. Thus, we set $\tilde{p}(o_{t})=\tilde{p}(o^{I}_{t})$. The self-prior is implemented using a normalizing-flows-based density estimator, specifically Neural Spline Flows (NSF) (Durkan et al.,, 2019). The agent autonomously learns the self-prior by maximizing the log-likelihood of the observation at each time step. For efficient training of neural networks, we sample observations from a replay buffer $\mathcal{D}$ that stores recently experienced trajectories:
Self-prior: $~{}~{}~{}~{}~{}\tilde{p}_{\xi}(o_{t})$ (13)
$\operatorname*{\arg\min}_{\xi}\mathcal{L_{\text{self}}}$ $=\operatorname*{\arg\min}_{\xi}\mathbb{E}_{o\sim\mathcal{D}}[-\log{\tilde{p}_{\xi}(o)}]$
Unlike the discrete model, which plans actions by minimizing the sum of expected free energy up to a future time $t+N$, the continuous model cannot enumerate such possibilities. To address this, we adopt a learning approach inspired by the actor-critic method in reinforcement learning, similar to (Millidge,, 2020). Specifically, we train a separate policy network that predicts an appropriate action $a_{t}$ from the state $s_{t}$, and additionally train an expected utility network that estimates expected free energy over an infinite horizon beyond $t+N$:
Policy: $q_{\theta}(a_{t}\mid s_{t})$ (14)
Expected utility: $g_{\psi}(s_{t})$
This approach is similar to Soft Actor-Critic (Haarnoja et al.,, 2018), where both the policy network and the utility network take the posterior feature $s_{t}$ as input, use a 3-layer MLP to output the parameters of a multivariate Gaussian distribution, and sample from this distribution. Accordingly, the policy and utility networks are trained as follows:
$\operatorname*{\arg\min}_{\theta}\mathcal{L}_{\text{policy}}$ $=\operatorname*{\arg\min}_{\theta}\sum\nolimits_{t}G^{\lambda}_{t}$ (15)
$\operatorname*{\arg\min}_{\psi}\mathcal{L}_{\text{utility}}$ $=\operatorname*{\arg\min}_{\psi}\sum\nolimits_{t}(g_{\psi}(s_{t})-G^{\lambda}_{t})$
$G^{\lambda}_{t}$ $=\mathcal{G}(s_{t})+\gamma_{t}\begin{cases}(1-\lambda)g_{\psi}(s_{t+1})+\lambda G^{\lambda}_{t+1},&\text{if}\ t<H\\ g_{\psi}(s_{H}),&\text{if}\ t=H\end{cases}$
We use GAE($\lambda$) estimation to train the utility network (Schulman et al.,, 2015). For all other unspecified hyperparameters and training stabilization techniques throughout the continuous model (e.g., KL-divergence clipping, action entropy scaling), we follow (Mazzaglia et al.,, 2021).
4 Experiments and Results
In this section, we validate the emergence of goal-directed behavior driven by the self-prior, which is learned from multimodal sensory experiences, through computer simulations. Our experimental setup assumes a scenario in which an infant agent detects a sticker on its left arm, pays attention to the object, and eventually removes it with its right hand—similar to the study by (Bigelow,, 1986), where a silent toy was placed on the body of a blind infant to examine reaching behavior.
Our experiment is based on the following assumptions: (i) since an infant typically does not have stickers attached to its body, it acquires a self-prior that "there are no stickers on my body"; (ii) when a sticker is attached, the mismatch between sensory observations and the self-prior increases free energy; and (iii) to reduce free energy, the agent spontaneously generates reaching behavior aimed at the sticker.
Through these experiments, we (i) verify the process by which the self-prior, representing the agent’s own preference, is autonomously formed, (ii) confirm the emergence of reaching behavior within the free energy minimization framework incorporating the self-prior, and (iii) in the discrete environment, show that behavior can vary for the same stimulus depending on changes in the self-prior.
4.1 Discrete Simulation
In the discrete environment, the agent receives multimodal observations composed of discrete proprioceptive and tactile inputs (Figure 3). The agent’s right hand can occupy one of five positions, labeled $0,1,2,3,$ or $4$, representing its proprioceptive state. The left arm has three tactile sensors, located at positions $1,2,$ and $3$ in the right hand’s coordinate frame, and each sensor can be activated independently, yielding $82^{3}=$ 8 possible tactile states. Consequently, the observation variable $o_{t}$ can take $5\times 8=40$ distinct values and is represented as a 40-dimensional one-hot column vector (Table 1).
Figure 3: Overview of the discrete environment. The right hand can move left or right either above the left arm or outside of it, and tactile input occurs either where the right hand is located or where the sticker is attached.
Table 1: Possible combinations of sensory observations in the discrete environment
$o_{t}$ Touch Hand position
(000 $\sim$ 111) (0 $\sim$ 4)
0 000 0
1 000 1
$\vdots$ $\vdots$ $\vdots$
28 101 3
$\vdots$ $\vdots$ $\vdots$
39 111 4
A caregiver can attach a sticker to the agent’s left arm. If a sticker is placed at a specific position (e.g., position $x$), the agent continuously receives a tactile signal from position $x$, regardless of the right hand’s position. If the right hand exactly matches the sticker’s location, the agent perceives a single tactile signal.
At each time step, the agent selects one of three actions $a_{t}$: (i) remain stationary, (ii) move one step to the left, or (iii) move one step to the right. Thus, $a_{t}$ is represented as a 3-dimensional one-hot column vector. In the discrete experiment, the policy $\pi$ considers all $813^{4}=$ 81 possible action sequences over a 4-step future horizon, and each sequence is probabilistically chosen according to the sum of expected free energy as defined in Equation 10.
In the first phase of the discrete-environment experiment, we let the agent operate without constraints for a sufficiently long period (10,000 steps). In the case of real infants, such observations may correspond to a variety of everyday experiences that fulfill extrinsic preferences. However, for simplicity, we assumed a motor babbling condition in which the agent randomly selects its actions $a$. Initially, the self-prior was uniformly distributed across all possible observations, so it did not respond even when a sticker was present (Figure 5(a)).
Through motor babbling, the agent learned the pattern that “a single tactile signal occurs when the right hand is over the left arm,” and as a result, “no tactile signal occurs when the right hand is off the left arm” (Left of Figure 4). When a caregiver attaches a sticker to position 1 on the left arm, a tactile signal is generated at a location different from the current hand position (position 4). According to the newly learned self-prior, this is a low-probability situation, resulting in an increase in free energy (i.e., attention is drawn due to the mismatch with the model). The agent then begins planning actions to reduce expected free energy and concludes that touching the sticker aligns best with its self-prior. As a result, the agent shows goal-directed movement toward position 1 (Figure 5(b)). Similarly, when the caregiver moves the sticker to position 3, the agent detects this new signal as incongruent with its model and promptly reaches toward position 3.
Figure 4: Change in self-prior over time. Before the sticker is attached, the probability increases for situations where no sticker is present on the arm ($t<10,000$). After the sticker is attached, the agent gradually adapts to the new situation where the sticker is present ($t\geq 10,000$).
(a) Before acquiring the self-prior ($t$ = 0)
(b) After acquiring the self-prior ($t$ = 10,000)
Figure 5: Comparison of the agent’s behavior before and after acquiring the self-prior. The top panel illustrates environmental changes over time: the red line indicates the hand position, the yellow dashed line indicates the sticker position. White areas denote where tactile feedback occurred, black areas indicate no tactile feedback, and gray areas represent regions outside the arm where tactile feedback never occurs. The bottom panel shows expected free energy over time. Each green dot represents the free energy of a candidate policy, with lower free energy policies being more likely to be selected. The red line connects the actually selected policies. (a) Before acquiring the self-prior, the agent does not respond even when a sticker is attached to the arm. (b) After acquiring the self-prior, goal-directed behavior emerges: the agent moves its hand to the sticker’s location when it appears on the arm.
Next, we let the agent undergo another round of motor babbling (20,000 steps) while the sticker remained at position 3. Over time, the self-prior was gradually updated so that the probability of tactile signal at position 3 increased relative to when the sticker was absent. In other words, the agent came to expect tactile input at position 3 regardless of its hand position (Right of Figure 4).
After this extended babbling, if the sticker was moved to position 1, the agent still noticed the mismatch (as a tactile signal at position 1 is inconsistent with the self-prior) and reached out to touch it. However, if the sticker was placed back at position 3, the agent no longer exhibited consistent goal-directed behavior (Figure 6). Because the self-prior had already adapted to position 3, no additional mismatch arose. Consequently, the difference in expected free energy between reaching or not was negligible, so the agent simply behaved randomly.
Figure 6: Agent’s behavior when a sticker has persistently been attached to position 3 ($t=30{,}000$). When the sticker is placed on other positions, the agent still reaches toward them; however, it no longer shows interest when the sticker is placed at position 3.
This situation, in which the self-prior does not induce any motivation, allows us to isolate the influence of the traditional information gain component of intrinsic motivation in the expected free energy framework. In this experiment, since the parameters $\mathbf{A}$ and $\mathbf{B}$ are already fixed in a fully accurate state, there is no new information to be gained from observations. As a result, no epistemic (information-seeking) drive emerges, and all policies yield nearly identical expected free energy, causing the agent to behave randomly.
In contrast, we observed that actions arising from the self-prior could be generated even when no additional information gain was possible. In other words, because the parameters of the self-prior, $\mathbf{C}$, continuously evolved with experience, intrinsic motivation related to familiarity based on experience was still induced. As a result, the agent was driven to act in ways that would align the newly updated self-prior with the observations.
4.2 Continuous Simulation
Up to this point, we have introduced the core idea that goal-directed behavior can emerge from the self-prior using a simple discrete setup. We now extend this to an environment with continuous-valued variables, inspired by the environment setup of (Marcel et al.,, 2022) (Figure 7).
[…]
To represent both tactile and proprioceptive modalities as a unified observation $o_{t}$ in the continuous model, we embed each modality into vectors of the same dimension. Specifically, we embed the tactile matrix using a four-layer Conv2D module and the proprioceptive vector using a two-layer MLP. The resulting embeddings are then combined via element-wise addition to yield the integrated observation embedding $o_{t}$. To decode from this embedding back to the original sensory modalities, we use a four-layer ConvTranspose2D for the tactile output and a two-layer MLP for the proprioceptive output.
[…]
Figure 7: Overview of the continuous environment. As in the discrete environment, the right hand can move over and around the left arm, and tactile sensations are generated where the hand or the sticker is located.
We conducted the continuous environment experiments on a machine running Ubuntu 22.04.5 LTS 64-bit (Linux 5.15.0-1066) with an Intel Xeon E5-2698 v4 CPU and an NVIDIA Tesla V100-SXM2 GPU, using Python 3.11.10 and PyTorch 2.5.1. For our normalizing-flows-based self-prior, we employed the “zuko” library (version 1.3.0), and we based much of our overall algorithmic implementation on Dreamer (Hafner et al.,, 2020) and Contrastive Active Inference (Mazzaglia et al.,, 2021).
To train and evaluate our deep model, we stored episodes in a replay buffer and randomly sampled them during model training. Each episode was collected either by using random actions or by following the agent’s policy from active inference, chosen with equal probability (50%). Likewise, for half of the episodes, a sticker was placed at a random location on the arm; for the other half, no sticker was used. Each episode consisted of $L=1000$ time steps.
After collecting an initial batch of 10 episodes, we trained the model for 100 epochs at the end of each episode. In each epoch, we sampled $B=50$ trajectories of length $L=50$ from the replay buffer, and generated imagined rollouts with a planning horizon $H=15$ for policy learning. For variational free energy minimization, trajectories were sampled from all available data—regardless of the presence of a sticker or whether the behavior was random or policy-driven. This was intended to ensure the acquisition of an accurate world model, similar to the assumption in the discrete environment that $\mathbf{A}$ and $\mathbf{B}$ were already known. Likewise, all data were used for expected free energy minimization, so that the agent could learn which actions reduce free energy effectively across various conditions.
For self-prior learning, we constructed the dataset such that only about 5% of the episodes included a sticker, while the rest were sampled without any sticker present. Consequently, as in the discrete environment experiment, the self-prior came to assign high probability to observations like “a single-point touch occurs when the right hand is above the left arm” and “no touch occurs when the right hand is away”.
(a) Time series of reaching behavior
(b) Change in expected free energy
Figure 8: Agent behavior when a sticker is placed in the continuous environment. (a) When a sticker (blue circle) is attached, the hand (red circle) moves toward it, illustrating goal-directed reaching behavior. The figure visualizes the agent’s tactile matrix, where grayscale tactile data is overlaid with colored markers for the hand and sticker for clarity. (b) Reaching for the sticker reduces expected free energy, and removing the sticker leads to its minimization. Shaded areas represent the standard deviation across 64 experiments using 8 model training seeds tested on 8 environment seeds. The figure uses seed 14, selected for clearly demonstrating the reaching behavior.
The agent displayed goal-directed reaching behavior whenever a sticker was placed on the left arm (Figure 8(a)). This occurred because touching the sticker minimized free energy by aligning with the self-prior (Figure 8(b)). Moreover, after successfully reaching the sticker, the agent persisted in touching it until the sticker was removed, indicating a consistent drive to resolve mismatch rather than merely losing interest once contact was made.
Interestingly, the agent exhibited reaching behavior toward the sticker even when the right hand was not initially touching the arm. If the self-prior had been learned based solely on tactile information, (i) the case where the left arm is touched by the right hand and (ii) the case where a sticker is attached might not be distinguishable—since both yield a single tactile point on the arm. In that case, no reaching behavior toward the sticker would emerge. However, the agent did reach toward the sticker, suggesting that the self-prior was learned by integrating both tactile and proprioceptive inputs. In this sense, our system, which utilizes a self-prior learned from multimodal sensory inputs to generate reaching behavior, resembles the concept of a body schema mentioned in (Gallagher,, 1986; Hoffmann,, 2021), which supports action planning and control.
Additionally, even if extra touches arose mid-trajectory—e.g., from the right hand itself—thus momentarily increasing mismatch relative to a single-point self-prior, the agent accepted this short-term deviation to achieve the ultimate goal of minimizing overall free energy. This robust, noise-tolerant sequence of goal-directed action demonstrates that our agent maintained a coherent intention throughout its interactions.
5 Discussion
We proposed the self-prior as an intrinsic preference that is autonomously formed from an agent’s sensory experiences. We then applied it to a simulated infant agent within the active inference framework and confirmed the emergence of goal-directed behavior—specifically, reaching for a sticker. In this section, to clarify the contributions of our work to developmental science, we first provide a detailed discussion of how the self-prior can be classified within existing computational models of intrinsic motivation. We then propose the role of the self-prior as a potential origin of early intentional behavior. After that, we explain how our approach differs from previous studies that attempted to generate spontaneous behaviors, and finally, we discuss the limitations of this work.
5.1 Internal Homeostatic Intrinsic Motivation
The self-prior represents an intrinsic motivation term that has been separated from extrinsic motivation, which may seem counterintuitive. Furthermore, the idea that intrinsic motivation can emerge from sensory data originating from the external world might appear paradoxical. We argue that this confusion stems from the inconsistent use of the terms “external” and “internal,” as also noted by (Oudeyer & Kaplan,, 2009). In this paper, we adopt the classification they recommend, distinguishing between:
  • External vs. internal, referring to whether the reward computation occurs in the environment or within the agent.
  • Extrinsic vs. intrinsic, referring to whether the criterion for reward is determined externally or internally.
According to these definitions, any external reward—such as a game score in machine learning tasks—qualifies as extrinsic, because it is provided from outside the agent, along with its preset threshold. Meanwhile, an internal drive can be either extrinsic or intrinsic. An internal drive is considered intrinsic if the agent calculates its own reward threshold based on its internal model—such as measuring “new information gain” according to a learned world model. By contrast, an internal yet extrinsic motivation, which can be confusing at first, is one in which the agent checks its own state and directly computes a reward based on a criterion that is externally fixed, for example “maintaining a battery level of 50%.”
Thus, the decisive factor for distinguishing extrinsic versus intrinsic motivation is not whether the information originates in the environment or within the agent, but rather who sets the threshold for reward computation. In this sense, the expected free energy remains an internal framework, since the agent itself computes both extrinsic and intrinsic components. Moreover, the self-prior approach introduced here qualifies as internal intrinsic motivation because the agent’s internal model (the self-prior) directly calculates and updates the criterion for behavior.
(Oudeyer & Kaplan,, 2009) also introduce another classification axis—homeostatic versus heterostatic—and note that homeostatic intrinsic motivation can indeed exist. In other words, if the agent’s drive is directed toward acquiring new information, it constitutes a heterostatic form of intrinsic motivation; however, maintaining a familiar state can also be considered an intrinsic motivation (homeostatic), because the agent internally defines the “familiarity” setpoint.
The expected free energy framework typically interprets the preferred prior as covering observations tied to survival requirements (e.g., blood sugar or battery level). Because such setpoints must not be altered arbitrarily, they are often considered extrinsic and not explicitly distinguished from their homeostatic nature. However, if the agent can freely specify its own setpoints, we can classify those setpoints as homeostatic intrinsic motivations.
Accordingly, we regard the motivation induced by the self-prior as an internal homeostatic intrinsic motivation—it aims to “maintain or restore a reference pattern” learned from the agent’s own sensory experiences (Table 2). This stands in contrast to the more traditional, heterostatic intrinsic motivations that seek new information for learning (e.g., reducing uncertainty about the hidden state). Instead, the self-prior directly triggers action to resolve any mismatch from the already acquired pattern of one’s own body. In (Oudeyer & Kaplan,, 2009), this is akin to distributional familiarity motivation (DFM), which drives the agent to return to sensory configurations it has repeatedly encountered.
Table 2: Positioning of the self-prior within a taxonomy of motivations
Extrinsic Intrinsic
Homeostatic Heterostatic
External Score in video game (N/A) (N/A)
Internal Gap of sensory data from fixed value Gap of sensory data from self-prior Information gain by sensory data
5.2 Origin of Early Intentional Behaviors
Infants initially depend heavily on reflex mechanisms for action. Over time, however, they gradually develop what (Mele & Moser,, 1994) describes as “intentional behavior,” comprising (i) a motivational system that judges how to handle incoming stimuli and (ii) the capability to execute genuinely goal-directed actions. According to (Zaadnoordijk & Bayne,, 2020), stimulus-driven intentions arise first, and then, with accumulating experience, more endogenous intentions emerge. For instance, newborns rely on reflexes such as rooting to acquire feeding experience; later, merely seeing a bottle prompts them to reach out and drink (stimulus-driven intention). As they become aware of internal states like hunger, they begin actively searching for the bottle themselves (endogenous intention).
Over time, to recognize themselves as independent entities, infants must undergo abundant experiences featuring multimodal redundancy and temporal–spatial contingencies (Rochat,, 1998). Sequentially developing intentional actions may spontaneously provide opportunities for such experiences. (Thelen et al.,, 1993) similarly noted that infants initially lack both the ability and the intention to perform reaching, and that the formation of stable intentions, coupled with active exploration driven by those intentions, facilitates the learning of reaching. However, prior studies have not examined in detail the neural architecture that underlies both stimulus-driven intentions and intentional actions.
Meanwhile, neurophysiological evidence suggests that the anterior cingulate cortex primarily deals with reward-related errors in individual decisions, whereas the posterior cingulate cortex processes discrepancy signals related to autobiographical information—such as one’s own body or past experiences—and triggers new behaviors (Pearson et al.,, 2011; Brewer et al.,, 2013). Our proposed self-prior aligns with this notion of “self-relevant predictions,” indicating that when new input fails to satisfy these predictions, the system initiates action. In other words, even without any externally designed goal, an intrinsic drive to preserve a self-prior learned from past experiences can manifest behaviors akin to an infant persistently touching a newly discovered sticker on their arm. This mirrors Piaget’s concept of primary circular reactions, where repeated attempts can refine intentions further, offering a theoretical foundation for computational studies on the origins of infant intentional behavior.
In our experiments, we fixed the self-prior’s learning rate. However, an excessively high rate could lead to overreactions (compulsive-like responses) to minor changes, whereas an overly low rate might cause indifference (apathy) toward new stimuli. Future work could investigate how more flexible learning rates—modeled after biological variability—shape different behavioral styles via self-relevant prediction-error processing.
5.3 Computation models for Intrinsic Motivation
Approaches that aim to computationally address intrinsic motivation have primarily been explored in reinforcement learning, where designing high-quality reward functions remains a significant challenge. As a result, intrinsic motivation has been proposed as a means for agents to more efficiently explore the environment and better infer reward functions. However, a fundamental question persists: how can an agent autonomously form motivation and generate behavior in environments where no goals are given, and what kind of value guides the agent’s actions under such conditions (Juechems & Summerfield,, 2019)?
An alternative approach is homeostatic reinforcement learning, in which the agent can generate a range of behaviors simply by maintaining certain sensory channels to specified thresholds, without environmental goal (Keramati & Gutkin,, 2011). Recent studies have extended this concept to complex settings using deep neural networks, examining how internal physiological states drive behavior (Yoshida et al.,, 2024). Under active inference, a similar mechanism emerges: by minimizing free energy with respect to an internal setpoint (expressed as a preferred prior), an agent can generate a single goal-directed act like reaching for a target (Oliver et al.,, 2022) or orchestrate repeated behaviors (Matsumoto et al.,, 2022).
However, although such methods may appear to operate without external rewards, most rely on the experimenter fixing the reference values (e.g., setpoints, preferred priors) in advance. Thus, while the agent’s reward calculation is internal, the ultimate source of its setpoint is extrinsic. As a result, these systems do not fully resolve the fundamental question of “Where does value come from, and how can the agent establish it autonomously?”
To allow the agent to establish its own value function independently, neither external rewards nor internal setpoints can be predetermined. But without these metrics, it is difficult to track progress or judge the utility of behaviors, which is why many studies on intrinsic motivation still use it primarily to facilitate exploration for extrinsic tasks (Aubret et al.,, 2023)—with limited focus on how the agent might self-generate and maintain goals.
Some works do investigate the behaviors arising purely from intrinsic motivation and qualitatively evaluate them in scenarios with no externally defined reward (Eysenbach et al.,, 2018; Pathak et al.,, 2017). While these highlight the agent’s acquisition of diverse skills or behaviors, they typically involve a heterostatic approach—continuously seeking novel states. Consequently, once uncertainty is sufficiently reduced, further exploration tends to diminish, and fewer new behaviors emerge.
Active inference offers a unified theoretical framework that can integrate both extrinsic and intrinsic motivations under the single quantitative measure of free energy (Parr et al.,, 2022). Furthermore, (Biehl et al.,, 2018) showed that many existing methods for intrinsic motivation can be cast within active inference as processes for improving the posterior—a perspective reminiscent of (Schmidhuber,, 2010), who defined intrinsic motivation in terms of seeking “better world models.” However, such posterior-improving motives are likewise heterostatic: they promote exploration until uncertainty is resolved, after which they elicit fewer notable actions.
Recently, studies have begun to explore homeostatic forms of intrinsic motivation, in which agents autonomously define their own setpoints but aim to preserve already learned states or past experiences, rather than continuously seeking novelty. For example, (Marcel et al.,, 2022) proposed a model in which a latent representation of self-touch is first learned and then used as a setpoint to guide the current observation toward the closest self-touch state, thereby inducing reaching behavior. Similarly, (Takemura et al.,, 2018; Kim et al.,, 2023) treated one-step forward and inverse models of a robot’s end-effector—learned from motor babbling—as behavioral setpoints, and demonstrated that such models could generate reaching toward distant targets without explicit rewards. Closer to our approach, (Sajid et al.,, 2021) introduced the concept of "learned preferences" within the free energy framework and semantically analyzed how a drive to reduce mismatches between observation and preference can lead to new behaviors. However, these approaches do not sufficiently discuss how such self-acquired setpoints could be integrated with conventional extrinsic or information-gain-based intrinsic motivations.
In this paper, we address these questions within the active inference framework by proposing a method whereby the agent shapes its own setpoint based on past experience, then spontaneously generates goal-directed behavior whenever incoming sensations deviate from this learned internal model. Our approach answers the question “Where does value come from?” by suggesting that “part of value may be shaped by the agent’s accumulated experience,” demonstrating that (i) the agent maintains self-defined setpoints instead of adhering to experimenter-imposed ones; (ii) it still utilizes the active inference machinery for long-horizon planning to minimize expected free energy; and (iii) it can be theoretically integrated with existing extrinsic or heterostatic intrinsic motivations. The result is a comprehensive framework wherein the agent’s value system is shaped and maintained internally, yet remains fully compatible with established motivational processes.
5.4 Limitations and Future Directions
Although our study provides a proof of concept, it is subject to several limitations. First, we tested our approach in a discrete environment and a low-degree-of-freedom (3-DoF) continuous environment. Real infants or robots, however, exhibit far more complex and varied sensorimotor interactions, underscoring the need to extend our framework to higher-dimensional state and action spaces. In particular, future work could draw on simulations of early human sensorimotor experience (including fetal stages) (Kim et al.,, 2022) to explore how rich, in-utero sensorimotor inputs shape self-priors and goal-directed actions, clarifying the types of actions that might emerge under embodied developmental constraints (Kuniyoshi,, 2019).
Additionally, our experiments focused on scenarios wherein the sticker is introduced abruptly. This choice helped isolate the effects of familiarity-based intrinsic motivation by excluding a fixed extrinsic motivation (such as externally imposed reward criteria). In reality, infants and robots must often balance multiple drives (e.g., energy homeostasis and novelty-driven information gain). Designing experiments that systematically assess how different motivations complement or conflict—and how these are integrated into coherent behavior—remains an important avenue for future research.
Our study demonstrated that, by separating an intrinsic homeostatic term based on a learnable self-prior from the extrinsic term based on a fixed preferred prior—commonly used in prior work with expected free energy—an agent can generate not only actions aimed at achieving predefined goals but also spontaneously generate goal-directed behaviors. However, a current limitation is that while the use of the self-prior is integrated into policy generation via expected free energy, the learning of the self-prior is performed separately from the learning of variational free energy.
One potential clue to resolving this separation lies in a hierarchical architecture. Specifically, a higher-level hidden state that reflects long-term experiences may be gradually learned, and the likelihood distribution inferred from it could serve as the self-prior for the lower level. This differs from conventional hierarchical inference, where the higher level simply learns priors and determines actions on a slower timescale to influence the lower level (e.g., $P(s_{t}|s_{t-1},a_{t-1})=\text{Cat}(\mathbf{B}_{a_{t-1}})$). Instead, it continuously receives information from the lower level to learn a prior, serving only to form a self-prior over a long duration (e.g., $P(s)=\text{Cat}(\mathbf{B})$).
Another possibility is that the learning of the self-prior is handled not by the cortex but by a separate memory system such as the hippocampus, and thus may require structural separation from the inference system proposed under the free energy principle. Concrete validation of these open hypotheses is left for future work.
In the discrete environment, we saw that placing a sticker and leaving it in place gradually altered the self-prior and, consequently, the agent’s behavior. By contrast, in the continuous environment, we fixed the probability of encountering a sticker at 5%, effectively decoupling self-prior learning from the evolving incidence of sticker observations. This occurred because the self-prior model learned from a replay buffer that stored past data without considering time in sequence, making it difficult to investigate incremental changes in the self-prior over time. To model how a stimulus-driven intention might become consolidated into a more endogenous form of intention in real infants or robots, we may need hierarchical state-space models capable of long-range temporal dependencies (Friston et al.,, 2017) or transformer-based architectures (Vaswani et al.,, 2017) that can learn world models from extended temporal sequences (Chen et al.,, 2022).
6 Conclusion
Building on the free energy principle and the active inference framework, we have demonstrated that an agent’s intrinsic drive to learn and maintain a “self-prior”—akin to a body schema—can induce goal-directed behaviors even in the absence of externally specified rewards. In particular, whereas the conventional active inference literature has emphasized heterostatic intrinsic motivation (i.e., seeking new information), our work highlights a homeostatic form of intrinsic motivation—one that actively strives to preserve familiar sensory experiences.
Using a simulated infant touching and examining a sticker on its arm as an illustrative example, we showed how an internally derived drive to resolve mismatches between the self model and incoming data can manifest in behaviors such as reaching and sticker removal, all without any explicit reward criteria. This offers a computational interpretation of stimulus-driven intentional actions in developmental psychology and suggests potential relevance to the neurobiological literature on posterior cingulate cortex, which processes errors linked to autobiographical information.
Future work should (i) incorporate both familiarity-based and additional motivations (extrinsic and information-seeking) to test how multiple drives may be integrated within a single free-energy framework, and (ii) extend our approach to longer time scales and more complex physical environments—such as high-DoF robotic platforms or full-scale infant simulations—to investigate the long-term formation and refinement of intentions. This will be particularly important for capturing the developmental trajectory of intentional behavior.
Acknowledgments and Disclosure of Funding
This work was supported by JST, PRESTO Grant Number JPMJPR23S4, Japan.