### v2

# Emergence of Goal-Directed Behaviors via Active Inference with Self-Prior

Abstract
Infants often exhibit goal-directed behaviors, such as reaching for a sensory stimulus, even when no external reward criterion is provided. These intrinsically motivated behaviors facilitate spontaneous exploration and learning of the body and environment during early developmental stages. Although computational modeling can offer insight into the mechanisms underlying such behaviors, many existing studies on intrinsic motivation focus primarily on how exploration contributes to acquiring external rewards. In this paper, we propose a novel density model for an agent’s own multimodal sensory experiences, called the “self-prior,” and investigate whether it can autonomously induce goal-directed behavior. Integrated within an active inference framework based on the free energy principle, the self-prior generates behavioral references purely from an intrinsic process that minimizes mismatches between average past sensory experiences and current observations. This mechanism is also analogous to the acquisition and utilization of a body schema through continuous interaction with the environment. We examine this approach in a simulated environment and confirm that the agent spontaneously reaches toward a tactile stimulus. Our study implements intrinsically motivated behavior shaped by the agent’s own sensory experiences, demonstrating the spontaneous emergence of intentional behavior during early development. 00footnotetext: * Co-corresponding authors: Hoshinori Kanazawa, Yasuo Kuniyoshi 00footnotetext: Code is available at https://github.com/kim135797531/self-prior.
Research Highlights
– Suggests a computational model for early intentional behavior, integrating body-schema formation and goal-directed actions under the free energy principle.
– Introduces a self-prior as an internal density model that drives the emergence of goal-directed behaviors without external rewards.
– Demonstrates spontaneous reaching for a sticker by minimizing mismatches between observed multimodal sensory inputs and the empirically acquired self-prior.
1 Introduction
Infants spontaneously exhibit goal-directed behaviors, such as reaching for a sensory stimulus or actively exploring their surroundings, even in the absence of external rewards essential for survival. These behaviors are highly motivated by internal satisfaction and are referred to as intrinsically motivated behaviors (czikszentmihalyi_flow_1990; ryan_intrinsic_2000). Such behaviors facilitate unsupervised learning without explicit external rewards or punishments, which is known to offer various advantages in early developmental stages (white_motivation_1959; gopnik_philosophical_2009; zaadnoordijk_lessons_2022; kanazawa_open-ended_2023). For example, sensory experiences arising from an infant’s spontaneous movements, such as self-touch, not only promote learning about one’s own body (e.g., acquiring body representations) (hoffmann_development_2017), but also contribute to the formation of an early sense of self (rochat_self-perception_1998).
Traditional observational studies and neuroscience methodologies have been employed to investigate the mechanisms underlying intrinsically motivated behavioral development (di_domenico_emerging_2017). In particular, computational modeling has also been utilized to quantitatively interpret and predict behavioral development (oudeyer_what_2009; shultz_computational_2013). Experiments using computer simulations allow researchers to freely manipulate and control specific variables to investigate various scenarios, making them advantageous for uncovering latent effects and interactions that may be difficult to observe through conventional behavioral studies alone.
Although computational modeling is a useful approach for studying infant behavioral development, many models proposed in modern robotics and machine learning tend to treat intrinsic motivation primarily as an exploratory mechanism for improving reward acquisition efficiency in environments with sparse external rewards (aubret_information-theoretic_2023). Indeed, such approaches can enhance reinforcement learning agents’ ability to improve explicit task performance, such as maximizing video game scores. However, they do not sufficiently address how entirely new behaviors emerge when no explicit external reward objective is set, meaning that no performance criterion or information gain is provided.
To address this limitation, we propose a novel computational model of intrinsic motivation based on the expected free energy (friston_free-energy_2010; friston_active_2016) (Figure 1). We introduce an internal density model, termed the self-prior, which enables an agent to learn a statistical representation of incoming multimodal sensory signals. When a mismatch between the learned probabilistic model and actual observations is detected, goal-directed behavior emerges without any explicit reward criterion.
Our approach combines several key features that, to our knowledge, have not been integrated in previous computational models of developmental behavior: (i) the behavioral setpoint is autonomously formed from the agent’s own experience rather than externally predefined; (ii) specific goal-directed behaviors such as reaching emerge independently, rather than serving merely as auxiliary mechanisms for reward exploration; (iii) behaviors persist even without information gain, as long as there is a mismatch with familiar states; and (iv) the mechanism is theoretically integrable with existing extrinsic and intrinsic motivations within the active inference framework. We implement a simulated agent that perceives a stimulus (such as a sticker) on its arm and spontaneously exhibits reaching behavior toward it through the self-prior mechanism. In the latter part of this paper, we discuss how the proposed mechanism may contribute to developmental science, positioning this study as a computational investigation into the early development of intentional behaviors in infants.
Figure 1: Emergence of reaching behavior via the self-prior and active inference. (a) When a sticker is placed on the left arm of the simulated agent, it detects a mismatch with its prior experience of not having a sticker, and reaches toward the sticker with its right hand to minimize the discrepancy. (b) Development of the self-prior through experience: as sensory experiences are collected, the probability distribution over sensory patterns gradually develops. (c) The active inference process in which the agent plans future actions to minimize expected free energy by aligning sensory inputs with the learned self-prior. As a result, the agent performs a reaching action toward the sticker. A full-body infant illustration is used for clarity, the actual experiment was conducted in a pseudo-3D environment.
1.1 Free Energy Principle
The free energy principle posits that an agent can learn to perceive and act by minimizing the "surprise" induced by sensory inputs (friston_free-energy_2010). This surprise refers to the prediction error that arises when the agent encounters unexpected sensory observations, and reducing this surprise becomes the agent’s primary objective.
At each time step $t$, the agent receives information from the world through its sensory organs. We denote this sensory input by $o_{t}$, referring to it as the observation. Although $o_{t}$ is all the agent can directly observe, there may exist underlying factors in the world that generate these observations. Such factors are called hidden states, denoted by $s_{t}$ at time $t$. For example, given a properly learned model, an agent can infer from a tactile pattern on its arm (observation $o_{t}$) that there is a sticker attached to its arm (hidden state $s_{t}$).
Consider an agent that must satisfy certain biological criteria for survival in a stable manner. To do so, the agent must accurately understand its environment, which amounts to inferring $p(o_{t})$. According to the free energy principle, an agent maintains its own probabilistic model of these hidden states because doing so enables it to approximate $p(o_{t})$ and thus make inferences about the world. In practice, we often maximize the log of $p(o_{t})$ (i.e., $\log p(o_{t})$) for computational convenience, which is referred to as maximizing the model evidence. Minimizing the negative term, $\log p(o_{t})$, known as surprise, is therefore equivalent to maximizing this model evidence.
In reality, calculating $p(o_{t})$ requires marginalization over hidden states (i.e., $p(o_{t},s_{t})$), which is often intractable. The variational free energy principle tackles this by introducing a variational approximation of Bayesian inference, wherein the agent indirectly reduces surprise by minimizing the variational free energy $\mathcal{F}$, an upper bound on $\log p(o_{t})$. Formally, $\mathcal{F}$ is defined as follows (see (parr_active_2022) for details):
$-\log p(o_{t})\leq\mathcal{F}$ $=\mathbb{E}_{q(s_{t})}[\log q(s_{t})-\log p(o_{t},s_{t})]$ (1)
$=\underbrace{D_{\mathrm{KL}}[q(s_{t})\|p(s_{t})]}_{\text{complexity}}-\underbrace{\mathbb{E}_{q(s_{t})}[\log p(o_{t}\mid s_{t})]}_{\text{accuracy}}$
where $D_{\mathrm{KL}}(\cdot\|\cdot)$ denotes the Kullback–Leibler divergence, and $q(\cdot)$ represents the variationally approximated distribution. That is, $q(s_{t})$ is the variational distribution that approximates the posterior distribution over hidden states. Therefore, to reduce surprise at time $t$, the agent must keep the posterior and prior distributions close (reducing complexity) while maximizing the likelihood of observations (improving accuracy).
1.2 Active Inference and Expected Free Energy
While the free energy principle describes how an agent updates its model to reduce surprise based on current sensory observations, the agent can also directly change its observations through actions $a_{t}$ to make them align better with its internal model. This is the essence of active inference, and more concretely, the approximate posterior $q(s_{t})$ is extended to include actions as $q(s_{t},a_{t})=q(a_{t}\mid s_{t})\,q(s_{t})$.
To obtain optimal actions $a_{t}$ for both present and future steps, a method has been proposed that computes expected free energy by considering the expected free energy at future time points (friston_active_2016). That is, the agent predicts multiple possible future states and selects the trajectory that is expected to minimize free energy most effectively.
However, since future observations cannot be obtained before the corresponding time actually arrives, the agent cannot directly compute the free energy of future states. Therefore, the expected free energy assumes that the agent possesses a preferred prior $\tilde{p}(o_{t})$ over desired observations, which acts as a setpoint. As a result, $p(o_{t},s_{t})$ is extended to include actions and is biased toward preferred observations, expressed as $\tilde{p}(o_{t},s_{t},a_{t})=p(a_{t})\,p(s_{t}\mid o_{t})\,\tilde{p}(o_{t})$. Thus, the expected free energy $\mathcal{G}$ is defined as follows (see (millidge_relationship_2020; mazzaglia_contrastive_2021) for details):
$\mathcal{G}$ $=\mathbb{E}_{q(o_{t},s_{t},a_{t})}[\log q(s_{t},a_{t})-\log\tilde{p}(o_{t},s_{t},a_{t})]$ (2)
$\approx-\underbrace{\mathbb{E}_{q(o_{t})}[\log\tilde{p}(o_{t})]}_{\text{extrinsic value}}-\underbrace{\mathbb{E}_{q(o_{t})}[D_{\mathrm{KL}}[q(s_{t}\mid o_{t})\|q(s_{t})]]}_{\text{intrinsic value}}-\underbrace{\mathbb{E}_{q(s_{t})}[\mathcal{H}(q(a_{t}\mid s_{t}))]}_{\text{action entropy}}$
where $\mathcal{H}$ denotes the information entropy. That is, to reduce surprise at a future time $t$, the agent must (i) increase the probability of preferred observations, (ii) obtain high information gain from those observations, and (iii) maintain diverse actions. As shown above, expected free energy integrates both pragmatic (extrinsic) and epistemic (intrinsic) value into a single expression, thereby naturally resolving the exploration–exploitation dilemma.
2 Self-Prior: Acquired Preference to Induce Goal-Directed Behavior
Under the expected free energy, agents do not require scalar rewards to be explicitly defined by the environment. This is because the agent is assumed to possess information about preferred observations as a preferred prior distribution, which serves as the setpoint for action planning.
However, since the preferred prior is typically given in a fixed form, the removal of external reward signals does not imply that the agent has autonomously generated its goal-directed behavior. Therefore, in this study, we allow the agent to autonomously determine the setpoint for its action planning by additionally learning a density model, which we refer to as the self-prior.
The self-prior represents "the probability density over the frequency of observed sensations that the agent has experienced," and is defined as a variable preferred prior that induces behavioral tendencies for the agent to maintain or re-experience familiar states. Unlike fixed preferred priors, this means that models with the same structure can develop different priors depending on different experiences, leading to the emergence of different behaviors. To provide an intuitive understanding of how the self-prior operates, this paper presents the example of spontaneous reaching toward and removal of a sticker attached to one’s own body. This is similar to the experimental setup of (bigelow_development_1986), where a silent toy was placed on the body of a blind infant to examine reaching behavior. Here, we interpret the reason why the infant agent removes a sticker attached to its body as being due to a mismatch between the prior of "oneself without a sticker" and "the currently observed self." Moreover, this prior has been autonomously and gradually developed by the agent through everyday experiences (in which no sticker was attached). Since this self-prior does not contain specific instructions to remove the sticker, it is distinguished from approaches that directly command reaching by providing fixed setpoints.
The implementation of the self-prior in the active inference framework is derived from a simple modification to the standard expected free energy formulation, separating the self-prior term from the original preferred prior term. Specifically, this is achieved by redefining each preferred observation $o$ in terms of two components: (i) an extrinsic observation $o^{E}$ that must be maintained to satisfy survival requirements, and (ii) an intrinsic observation $o^{I}$ related to the self-prior. Here, $o^{I}$ refers to bodily states or sensations such as a tactile pattern on the arm, that can vary without threatening the agent’s essential functions. Consequently, we decompose the preferred prior $\tilde{p}(o)$ into $\tilde{p}(o^{E},o^{I})$, and in this study, we assume independence so that $\tilde{p}(o^{E},o^{I})\approx\tilde{p}(o^{E})\,\tilde{p}(o^{I})$:
Preferred prior: $\tilde{p}(o)=\tilde{p}(o^{E},o^{I})\approx\tilde{p}(o^{E})\,\underbrace{\tilde{p}(o^{I})}_{\mathclap{\textbf{Self-Prior}}}$~~~ (3)
In our approach, we choose to define the self-prior over observations $o$ rather than internal states $s$. This is because, as mentioned earlier, the self-prior can be derived through simple manipulation of the standard expected free energy formulation, enabling natural integration of fixed extrinsic preferences and experience based intrinsic preferences within the same framework. The fact that the self-prior is related to intrinsic motivation yet can emerge from external sensory data may seem counterintuitive. This confusion stems from inconsistent use of "external-extrinsic" and "internal-intrinsic" terminology. To resolve this confusion and rigorously classify intrinsic motivation, we adopt the computational classification of intrinsic motivation proposed by (oudeyer_what_2009):
  • External vs. Internal: Referring to whether the reward computation occurs in the environment or within the agent.
  • Extrinsic vs. Intrinsic: Referring to whether the reward criterion is determined externally or internally.
According to these definitions, any reward provided from the external environment, such as a game score, qualifies as extrinsic motivation. Meanwhile, internal motivation can be either extrinsic or intrinsic. An internal motivation is considered intrinsic if the agent sets its reward criterion according to its internal model (e.g., measuring information gain based on a learned world model). By contrast, an internal yet extrinsic motivation is one in which the agent checks its internal state and computes a reward based on an externally fixed criterion (e.g., "maintaining a battery level of 50%").
Therefore, the key to distinguishing extrinsic and intrinsic motivation is not the source of information, but rather who sets the reward criterion. In this sense, expected free energy is an internal framework since the agent itself computes both extrinsic and intrinsic components. Furthermore, the self-prior proposed in this paper qualifies as internal intrinsic motivation because the agent’s internal model directly calculates and updates the reward criterion.
Additionally, (oudeyer_what_2009) proposed that motivation can also be distinguished as homeostatic vs. heterostatic. For example, if there is an intrinsic motivation to pursue new information acquisition, this is heterostatic, whereas if there is an intrinsic motivation to maintain a familiar state, this is defined as homeostatic.
In active inference, the preferred prior has traditionally been interpreted as a fixed setpoint representing observations essential for survival (e.g., blood sugar level, battery level). These setpoints have been considered extrinsic values since they are externally specified, but we regard this term as strictly representing a homeostatic value characteristic. Therefore, the decomposition in Equation (3) separates the homeostatic value term in Equation (2) into extrinsic and intrinsic terms:
$-\underbrace{\mathbb{E}_{q(o_{t})}[\log\tilde{p}(o_{t})]}_{\text{homeostatic value}}=-\underbrace{\mathbb{E}_{q(o^{E}_{t})}[\log\tilde{p}(o^{E}_{t})]}_{\text{extrinsic (homeostatic)}}-\underbrace{\mathbb{E}_{q(o^{I}_{t})}[\log\tilde{p}(o^{I}_{t})]}_{\text{intrinsic (homeostatic)}}$ (4)
Note that unlike Equation (2), where $\mathbb{E}_{q(o_{t})}[\log\tilde{p}(o_{t})]$ is labeled as extrinsic value, Equation (4) labels it as homeostatic. Thus, the expected free energy incorporating the self-prior becomes:
$\mathcal{G}$ $\approx-\underbrace{\mathbb{E}_{q(o^{E}_{t})}[\log\tilde{p}(o^{E}_{t})]}_{\text{extrinsic (homeo-, fixed)}}-\underbrace{\mathbb{E}_{q(o^{I}_{t})}[\log\tilde{p}(o^{I}_{t})]}_{\text{intrinsic (homeo-, familiarity)}}-\underbrace{\mathbb{E}_{q(o_{t})}[D_{\mathrm{KL}}[q(s_{t}|o_{t})\|q(s_{t})]]}_{\text{intrinsic (hetero-, novelty)}}$-$\mathcal{H^{\prime}}$ (5)
Here, $\mathcal{H}^{\prime}$ denotes the entropy of the action distribution.
In summary, the self-prior in this study aims to "maintain or restore a reference pattern" learned from the agent’s sensory experiences, thus qualifying as internal homeostatic intrinsic motivation. This most closely resembles distributional familiarity motivation (DFM) among the classifications of intrinsic motivation proposed by (oudeyer_what_2009). Table 1 summarizes the positioning of the self-prior within this motivational taxonomy.
Table 1: Positioning of the self-prior within a taxonomy of motivations
Extrinsic Intrinsic
Homeostatic Heterostatic
External Score in video game (N/A) (N/A)
Internal
Gap of sensory data
from fixed value
Gap of sensory data
from self-prior
Information gain
by sensory data
As an alternative approach to implementing the self-prior concept, there also exist approaches that construct preferences over internal states. This is the approach proposed by (sajid_exploration_2021), but it could not explicitly integrate extrinsic and intrinsic motivations. Another approach could consider hierarchical models wherein internal representations learned at higher levels are transformed into self-priors for lower levels. We revisit these potential future extensions in the Discussion section.
3 Computational Model of Self-Prior
In this section, we introduce the concrete implementation of our model for computational simulations based on the active inference framework incorporating the self-prior. The overall framework is constructed using small-sized matrices in the discrete environment to allow direct inspection of model behavior, and deep neural networks using techniques developed in deep reinforcement learning in the continuous environment.
We implement this as a general-purpose structure applicable to a wide range of behaviors beyond specific tasks, using an end-to-end learning model that requires no prior preparation such as filtering specific parts of sensory data for the self-prior. Specifically, in the discrete environment model, we learn maximum likelihood estimation via empirical frequency normalization (plug-in MLE), while in the continuous environment model, we train Normalizing Flow based density estimators via maximum likelihood (durkan_neural_2019).
3.1 Model for Discrete Environment
The model for the discrete environment uses categorical distributions as the generative model, handling discrete random variables:
Prior: $P(s_{t}\mid s_{t-1},a_{t-1})=\text{Cat}(\mathbf{B}_{a_{t-1}})$ (6)
Posterior: $Q(s_{t})=\text{Cat}(\phi_{t})$
Likelihood: $P(o_{t}\mid s_{t})=\text{Cat}(\mathbf{A})$
Here, uppercase letters P,QP,\,Q denote discrete distributions, and $\mathbf{A},\,\mathbf{B},\,\phi_{t}$ are parameter matrices and vectors of categorical distributions. $\mathbf{B}$ implements the transition model, with separate $\mathbf{B}_{a}$ matrices for each action $a$. Using this notation, minimizing variational free energy yields the optimal posterior distribution parameter $\phi_{t}$ for the current observation $o_{t}$ through the Softmax normalization function $\sigma$ as follows (derivation process see Appendix A.1.1):
$\phi_{t}$ $\approx\sigma(\log{(\mathbf{A}\cdot o_{t})}+\log{(\mathbf{B}_{a_{t-1}}\phi_{t-1})})$ (7)
We assume that the likelihood distribution $\text{Cat}(\mathbf{A})$ and the prior distribution $\text{Cat}(\mathbf{B})$ have already been learned accurately. That is, the parameter $\mathbf{A}$ perfectly knows which observation $o_{t}$ corresponds to each possible hidden state $s_{t}$, and $\mathbf{B}$ perfectly knows the outcome of each possible action $a_{t}$ for each state. This simplifying assumption is intended to exclude side effects that may arise from other parts of the generative model when investigating changes in behavior due to changes in the self-prior (the learning of both likelihood and prior distributions is included in the model for the continuous environment discussed later).
The distribution $\tilde{P}(o_{t})$ that serves as the setpoint for future observations is given in the form of a categorical distribution $\text{Cat}(\mathbf{C})$, where $\mathbf{C}$ is a column vector with a row for each possible $o_{t}$. According to the definition in Eq. (3), the preferred prior $\tilde{p}(o_{t})$ should decompose into an extrinsic component $\tilde{p}(o^{E}_{t})$ and an intrinsic component $\tilde{p}(o^{I}_{t})$, but in this study, we excluded the extrinsic component in order to investigate purely intrinsic motivational behavior driven by the self-prior. Therefore, all components of the observation vector $o_{t}$ correspond to intrinsic observations $o^{I}_{t}$, and the entire $\text{Cat}(\mathbf{C})$ is defined as the self-prior $\tilde{P}(o^{I}_{t})$:
Self-prior: $\tilde{P}(o^{I}_{t})=\text{Cat}(\mathbf{C})$ (8)
The agent autonomously learns the self-prior by recording the frequency of each observation experienced so far and storing the corresponding observation probabilities in the parameter $\mathbf{C}$. The initial value of $\mathbf{C}$ is set so that all observations have equal probability. To determine the optimal action $a_{t}$ based on this, we compute the expected free energy for each action (derivation process see Appendix A.1.2):
$\mathcal{G}\approx(\mathbf{B}_{a_{t}}\phi_{t})\cdot\mathcal{H}[\mathbf{A}]+D_{\mathrm{KL}}[\mathbf{A}\mathbf{B}_{a_{t}}\phi_{t}\|\mathbf{C}]$ (9)
To evaluate future policies, we propagate the current belief $\phi_{t}$ through the transition model $\mathbf{B}$ into future beliefs $\phi_{t+1:t+N}$ along "imagined time": $\phi_{t+\tau}=\mathbf{B}_{a_{t+\tau-1}}\phi_{t+\tau-1}$. The optimal policy is selected by summing expected free energies for each candidate policy $\pi=\{a_{t},...,a_{t+N}\}$. The policy is expressed as a categorical distribution, from which actions are sampled:
Policy: $P(\pi)=\sigma(-\mathcal{G})$ (10)
3.2 Model for Continuous Environment
In prior research attempting to apply the free energy principle to (high-dimensional) continuous environments, many implementations use joint angles directly as internal states and generate actions through analytically derived backpropagation (e.g., (sancaktar_end_2020; priorelli_deep_2023)). While these approaches offer the advantage of analytical computation, they are limited in that the representation is constrained to joint angles, making it difficult to implicitly encode complex sensorimotor information in high-dimensional latent spaces and challenging to apply to complex tasks requiring long-term planning.
In contrast, approaches utilizing deep neural networks do not directly mimic biological details, but they provide a practical method to extend the core computational principles of active inference to high-dimensional problems (millidge_deep_2020). The focus of this study is to propose and demonstrate the conceptual mechanism of the self-prior, which is not bound to any specific neural implementation. Therefore, we follow the deep neural network based approach to enable representation of complex sensorimotor information in high-dimensional latent spaces and to perform future planning. Specifically, in this study, we follow the approach of prior research that interprets the Recurrent State-Space Model (RSSM) in PlaNet (hafner_planet_2019) as optimizing variational free energy (catal_learning_2020), and interprets policy gradient methods as minimizing expected free energy (millidge_deep_2020). The RSSM is essentially an extension of the Variational Autoencoder (VAE, (kingma2013auto)) architecture to sequential data, directly implementing the principle of variational free energy minimization.
Whereas the discrete model fixed the parameters of the likelihood and prior distributions in variational free energy, the continuous model follows the standard RSSM structure and trains the parameters $\phi$ using deep neural networks. The RSSM separates the deterministic latent state $h_{t}$ and the stochastic latent state $s_{t}$, enabling proper sequential decomposition:
Deterministic state: $h_{t}=f_{\phi}(h_{t-1},s_{t-1},a_{t-1})$ (11)
Stochastic prior: $p_{\phi}(\hat{s}_{t}\mid h_{t})$
Stochastic posterior: $q_{\phi}(s_{t}\mid h_{t},o_{t})$
Likelihood: $p_{\phi}(o_{t}\mid h_{t},s_{t})$
The deterministic state captures temporal dependencies through a GRU cell, while the stochastic prior and posterior each generate Gaussian distribution parameters from the deterministic states (and observations). This separation ensures proper variational decomposition at each time step (implementation details in Appendix A.2.2).
(a) Original active inference
(b) Active inference with self-prior
Figure 2: Graphical model of active inference using deep neural networks that minimize variational free energy $\mathcal{F}$ and expected free energy $\mathcal{G}$. The self-prior $\tilde{p}(o^{I}_{t})$ is trained to maximize the log-likelihood of observations $o^{I}_{t}$. In the expected free energy calculation (highlighted in blue), the learned self-prior serves as the behavioral setpoint alongside the fixed preferred prior. Although the preferred prior and self-prior can theoretically be applied simultaneously, we use only the self-prior in this study for clarity of exposition; thus, the preferred prior is shown faded in the figure.
The continuous model also investigates purely intrinsic motivational behavior, thus $\tilde{p}(o_{t})=\tilde{p}(o^{I}_{t})$. The self-prior is learned using Neural Spline Flows (NSF, (durkan_neural_2019)) to maximize observation log-likelihood. This is done by sampling observations $o$ from a replay buffer $\mathcal{D}$ that stores recently experienced trajectories:
Self-prior: $~~~~~\tilde{p}_{\xi}(o_{t})$ (12)
$\operatorname*{\arg\min}_{\xi}\mathcal{L_{\text{self}}}$ $=\operatorname*{\arg\min}_{\xi}\mathbb{E}_{o\sim\mathcal{D}}[-\log{\tilde{p}_{\xi}(o)}]$
Since action candidates cannot be enumerated in continuous environments, we use policy gradient methods. Following (millidge_deep_2020), we train a policy network that predicts actions from states and a value network that estimates expected free energy over an infinite horizon:
Policy: $q_{\theta}(a_{t}\mid s_{t})$ (13)
Expected utility: $g_{\psi}(s_{t})$
This is a similar approach to Soft Actor-Critic (haarnoja_sac_2018), and approximates expected utility through GAE($\lambda$) estimation (schulman_gae_2015). Training details and hyperparameters are provided in Appendix A.2.3 and (mazzaglia_contrastive_2021).
4 Experiments and Results
In this section, we validate the emergence of goal-directed behavior driven by the self-prior, which is learned from multimodal sensory experiences, through computer simulations. As briefly mentioned earlier, our experiment is based on the following assumptions: (i) since an infant typically does not have stickers attached to its body, it acquires a self-prior that "there are no stickers on my body"; (ii) when a sticker is attached, the mismatch between sensory observations and the self-prior increases free energy; and (iii) to reduce free energy, the agent spontaneously generates reaching behavior aimed at the sticker. Importantly, we did not preset any preferred states such as removing stickers or adopting specific postures; all goal-directed behaviors emerge from the self-prior autonomously formed from the sensations the agent experiences through motor babbling.
Through these experiments, we (i) verify the process by which the self-prior, representing the agent’s own preference, is autonomously formed, (ii) confirm the emergence of reaching behavior within the free energy minimization framework incorporating the self-prior, and (iii) in experiments in the discrete environment, additionally show that behavior can vary for the same stimulus depending on changes in the self-prior.
4.1 Simulation in Discrete Environment
In the discrete environment, the agent receives multimodal observations composed of discrete proprioceptive and tactile inputs (Figure 3). The agent’s right hand can occupy one of five positions, labeled $0,1,2,3,$ or $4$, representing its proprioceptive state. The left arm has three tactile sensors, located at positions $1,2,$ and $3$ in the right hand’s coordinate frame, and each sensor can be activated independently, yielding $2^{3}=8$ possible tactile states. Consequently, the observation variable $o_{t}$ can take $5\times 8=40$ distinct values and is represented as a 40-dimensional one-hot column vector (Table 2).
Figure 3: Overview of the discrete environment. The right hand can move left or right either above the left arm or outside of it, and tactile input occurs either where the right hand is located or where the sticker is attached.
Table 2: Possible combinations of sensory observations in the discrete environment
$o_{t}$ Touch Hand position
(000 $\sim$ 111) (0 $\sim$ 4)
0 000 0
1 000 1
$\vdots$ $\vdots$ $\vdots$
28 101 3
$\vdots$ $\vdots$ $\vdots$
39 111 4
A caregiver can attach a sticker to the agent’s left arm. If a sticker is placed at a specific position (e.g., position $x$), the agent continuously receives a tactile signal from position $x$, regardless of the right hand’s position. If the right hand exactly matches the sticker’s location, the agent perceives a single tactile signal.
At each time step, the agent can choose one of three actions $a_{t}$: (i) remain stationary, (ii) move one step to the left, (iii) move one step to the right. Thus, $a_{t}$ is represented as a 3-dimensional one-hot column vector. In the discrete environment experiment, there are $3^{4}=81$ possible policies $\pi$, which are sequences of actions up to 4 steps into the future, and are probabilistically determined by Equation 10 according to the sum of expected free energies for each.
At the beginning of the experiment, the self-prior was uniformly initialized across all possible observations. This makes the expected free energy equal for all policies, causing random policy selection. That is, no special response occurs even when a sticker is attached (Figure 5(a)).
Subsequently, we let the agent perform motor babbling without constraints for a sufficiently long period (10,000 steps). In the case of real infants, such experiences may correspond to a variety of everyday experiences that fulfill extrinsic preferences. However, for simplicity, we performed motor babbling in which the agent randomly selects its actions $a$. Through this process, the pattern that "a single tactile signal occurs when the right hand is over the left arm" is learned, and as a result, "no tactile signal occurs when the right hand is off the left arm" (Left of Figure 4).
At this point, when a caregiver attaches a sticker to position 1 on the left arm, a tactile signal occurs at a location different from the current hand position (position 4). According to the newly learned self-prior, this is a low-probability situation, resulting in an increase in free energy (i.e., attention is drawn due to the mismatch with the model). Accordingly, the agent begins planning actions to reduce expected free energy and concludes that touching the sticker aligns best with its self-prior. As a result, the agent shows goal-directed movement toward position 1 (Figure 5(b)). Similarly, when the caregiver moves the sticker to position 3, the agent detects this new signal as incongruent with its model and immediately reaches toward position 3.
Figure 4: Change in self-prior over time. Before the sticker is attached, the probability increases for situations where no sticker is present on the arm ($t<10,000$). After the sticker is attached, the agent gradually adapts to the new situation where the sticker is present ($t\geq 10,000$).
(a) Before acquiring the self-prior ($t$ = 0)
(b) After acquiring the self-prior ($t$ = 10,000)
Figure 5: Comparison of the agent’s behavior before and after acquiring the self-prior. The top panel illustrates environmental changes over time: the red line indicates the hand position, the yellow dashed line indicates the sticker position. White areas denote where tactile feedback occurred, black areas indicate no tactile feedback, and gray areas represent regions outside the arm where tactile feedback never occurs. The bottom panel shows expected free energy over time. Each green dot represents the free energy of a candidate policy, with lower free energy policies being more likely to be selected. The red line connects the actually selected policies. (a) Before acquiring the self-prior, the agent does not respond even when a sticker is attached to the arm. (b) After acquiring the self-prior, goal-directed behavior emerges: the agent moves its hand to the sticker’s location when it appears on the arm.
Next, we let the agent undergo another round of motor babbling (20,000 steps) while the sticker remained at position 3. Over time, the self-prior was gradually updated so that the probability of tactile signal at position 3 increased relative to when the sticker was absent. In other words, the agent came to expect tactile input at position 3 regardless of its hand position (Right of Figure 4).
After this extended babbling, if the sticker was moved to position 1, the agent still noticed the mismatch (as a tactile signal at position 1 is inconsistent with the self-prior) and reached out to touch it. However, if the sticker was placed back at position 3, the agent no longer exhibited consistent goal-directed behavior (Figure 6). Because the self-prior had already adapted to position 3, no additional mismatch arose. Consequently, the difference in expected free energy between reaching or not was negligible, so the agent simply behaved randomly.
Figure 6: Agent’s behavior when a sticker has persistently been attached to position 3 (t=30,000t=30{,}000). When the sticker is placed on other positions, the agent still reaches toward them; however, it no longer shows interest when the sticker is placed at position 3.
This situation, in which the self-prior does not induce any motivation, allows us to isolate the influence of the traditional information gain component of intrinsic motivation in the expected free energy framework. In this experiment, since the parameters $\mathbf{A}$ and $\mathbf{B}$ are already fixed in a fully accurate state, there is no new information to be gained from observations. As a result, no epistemic (information-seeking) drive emerges, and all policies yield nearly identical expected free energy, causing the agent to behave randomly.
In contrast, we observed that actions arising from the self-prior could be generated even when no additional information gain was possible. In other words, because the parameters of the self-prior, $\mathbf{C}$, continuously evolved with experience, intrinsic motivation related to familiarity based on experience was still induced. As a result, the agent was driven to act in ways that would align the newly updated self-prior with the observations.
4.2 Simulation in Continuous Environment
Up to this point, we have introduced the core idea that goal-directed behavior can emerge from the self-prior using a simple discrete setup. We now extend this to an environment with continuous-valued variables, inspired by the environment setup of (marcel_learning_2022) (Figure 7).
Initially, the agent’s two arms lie on the $xy$-plane. The lengths of the left forearm, left arm, torso, right arm, and right forearm are $80$, $70$, $140$, $70$, and $80\,\text{mm}$, respectively. The left forearm features a $30\,\text{mm}$-wide two-dimensional tactile sensor array, placed laterally relative to the forearm itself. At the endpoint of the right forearm is a circular "right hand" of radius $6\,\text{mm}$, which can generate tactile input on the left forearm.
All joint angles are expressed in radians. The left elbow and left shoulder angles are fixed at $\pi/3$ and $2\pi/3$, while the right shoulder and right elbow angles have respective ranges of $0,\,2\pi/3]$ and $0,\,3\pi/4]$. Additionally, the right hand can move along a pseudo-$z$ (height) axis within the interval $0,\,20]\text{mm}$.
Similar to the discrete case, the agent perceives both proprioception and touch; however, these quantities are now continuous. The proprioceptive input is a three-dimensional real-valued vector $(\text{shoulder},\text{elbow},\text{hand})$, while tactile input is represented by an $80\times 30$ real-valued matrix normalized to [0, 1][0,\,1]. Whenever the right hand is above the left forearm at a height of $\leq 10\,\text{mm}$, the tactile sensor on the left forearm is activated with intensity inversely proportional to the hand’s height.
To represent both tactile and proprioceptive modalities as a unified observation $o_{t}$ in the model for the continuous environment, we convert each modality into an embedding of the same dimensionality and integrate them. The tactile matrix is transformed into an embedding using a CNN, and the proprioceptive vector is transformed into an embedding using a MLP. We combine the two embeddings via element-wise addition to yield the final integrated observation embedding $o_{t}$. To decode from the embedding back to the original sensations, we use inverse transformation modules (see Appendix A.2.1 for detailed hyperparameters).
A caregiver can attach a sticker of radius $4\,\text{mm}$ to the left forearm at a fixed height of $0\,\text{mm}$. As a simplified environmental rule to demonstrate the core self-prior mechanism, the agent can remove the sticker if the center of its right hand remains within the sum of the hand’s and sticker’s radii for $10$ consecutive time steps. This represents a simplified assumption wherein sustained contact between the hand and sticker leads to object removal, without modeling complex manipulation actions such as grasping or pushing.
At every time step, the right shoulder and right elbow angles can each rotate between 0.05-0.05 and $0.05\,\text{rad}$, and the hand’s height can shift by 0.5-0.5 to $0.5\,\text{mm}$. To keep the hand near the forearm, we impose an additional constraint that the hand’s center must remain within $15\,\text{mm}$ of the forearm’s surface. If a proposed movement would place the hand outside this region, the agent’s action is disregarded and its previous position is retained.
Figure 7: Overview of the continuous environment. As in the discrete environment, the right hand can move over and around the left arm, and tactile sensations are generated where the hand or the sticker is located.
We conducted the continuous environment experiments on a machine running Ubuntu 22.04.5 LTS 64-bit (Linux 5.15.0-1066) with an Intel Xeon E5-2698 v4 CPU and an NVIDIA Tesla V100-SXM2 GPU, using Python 3.11.10 and PyTorch 2.5.1. For our normalizing flows based self-prior, we employed the “zuko” library (version 1.3.0), and we based much of our overall algorithmic implementation on Dreamer (hafner_dreamer_2020) and Contrastive Active Inference (mazzaglia_contrastive_2021).
To train and evaluate our deep model, we stored episodes in a replay buffer and randomly sampled them during model training. Each episode was collected either by using random actions or by following the agent’s policy from active inference, chosen with equal probability (50%). Likewise, for half of the episodes, a sticker was placed at a random location on the arm; for the other half, no sticker was used.
After collecting an initial batch of 100 episodes, we trained the model for 100 epochs at the end of each episode. In each epoch, we sampled $B=50$ trajectories of length $L=50$ from the replay buffer, and generated imagined rollouts with a planning horizon $H=15$ for policy learning. For variational free energy minimization, trajectories were sampled from all available data, regardless of the presence of a sticker or whether the behavior was random or policy-driven. This was intended to ensure the acquisition of an accurate world model, similar to the assumption in the discrete environment that $\mathbf{A}$ and $\mathbf{B}$ were already known. Likewise, all data were used for expected free energy minimization, so that the agent could learn which actions reduce free energy effectively across various conditions.
For self-prior learning, we constructed the dataset such that only about 5% of the episodes included a sticker, while the rest were sampled without any sticker present. Consequently, as in the discrete environment experiment, the self-prior came to assign high probability to observations like “a single-point touch occurs when the right hand is above the left arm” and “no touch occurs when the right hand is away”.
After collecting and training for 2,000 episodes, we attached a sticker to the agent’s left arm. As a result, the agent exhibited goal-directed behavior of reaching toward the sticker (Figure 8(a)). This occurred because touching the sticker aligned with the self-prior, minimizing free energy (Figure 8(b)). Moreover, after reaching the sticker, the agent continued touching it and eventually removed it, indicating that it had a persistent motivation to reduce the mismatch rather than simply losing interest after contact.
(a) Time series of reaching behavior
(b) Change in expected free energy
Figure 8: Agent behavior when a sticker is placed in the continuous environment. (a) When a sticker (blue circle) is attached, the hand (red circle) moves toward it, illustrating goal-directed reaching behavior. The figure visualizes the agent’s tactile matrix, where grayscale tactile data is overlaid with colored markers for the hand and sticker for clarity. (b) Reaching for the sticker reduces expected free energy, and removing the sticker leads to its minimization. Shaded areas represent the standard deviation across 64 experiments using 8 model training seeds (0˜7) tested on 8 environment seeds (0˜7). The figure uses seed 4, selected for clearly demonstrating the reaching behavior.
Interestingly, the agent exhibited reaching behavior toward the sticker even when the right hand was not initially touching the arm. If the self-prior had been learned based solely on tactile information, (i) the case where the left arm is touched by the right hand and (ii) the case where a sticker is attached might not be distinguishable, since both yield a single tactile point on the arm. In that case, no reaching behavior toward the sticker would emerge. However, the agent did reach toward the sticker, suggesting that the self-prior was learned by integrating both tactile and proprioceptive inputs. In this sense, our system, which utilizes a self-prior learned from multimodal sensory inputs to generate reaching behavior, resembles the concept of a body schema mentioned in (gallagher_body_1986; hoffmann_body_2021), which supports action planning and control.
Additionally, even if extra touches arose mid-trajectory (e.g., from the right hand itself) thus momentarily increasing mismatch relative to a single-point self-prior, the agent accepted this short-term deviation to achieve the ultimate goal of minimizing overall free energy. This robust, noise-tolerant sequence of goal-directed action demonstrates that our agent maintained a coherent intention throughout its interactions.
5 Discussion
We proposed the self-prior as an intrinsic preference that is autonomously formed from an agent’s sensory experiences. We then applied it to a simulated infant agent within the active inference framework and confirmed the emergence of goal-directed behavior, specifically reaching for a sticker. In this section, to clarify the contributions of our work to developmental science, we explain how our approach differs from previous studies that attempted to generate spontaneous behaviors, propose the role of the self-prior as a potential origin of early intentional behavior, and finally discuss the limitations of this work and future research directions.
5.1 Computation Models of Intrinsic Motivation
Approaches that aim to computationally address intrinsic motivation have primarily been explored in reinforcement learning, where designing high-quality reward functions remains a significant challenge. As a result, intrinsic motivation has been proposed as a means for agents to more efficiently explore the environment and better infer reward functions. However, a fundamental question persists: how can an agent autonomously form motivation and generate behavior in environments where no goals are given, and what kind of value guides the agent’s actions under such conditions (juechems_where_2019)?
Extrinsic Homeostatic Reinforcement Learning
An alternative approach is homeostatic reinforcement learning, in which the agent can generate a range of behaviors simply by maintaining certain sensory channels to specified thresholds, without environmental goal (keramati_reinforcement_2011). Recent studies have extended this concept to complex settings using deep neural networks, examining how internal physiological states drive behavior (yoshida_emergence_2024). Under active inference, a similar mechanism emerges: by minimizing free energy with respect to an internal setpoint (expressed as a preferred prior), an agent can generate a single goal-directed act like reaching for a target (oliver_empirical_2022) or orchestrate repeated behaviors (matsumoto_goal-directed_2022).
However, although these approaches appear intrinsic because behaviors occur without external rewards, most have a limitation in that the experimenter fixes the reference values (e.g., setpoints or preferred priors) in advance. That is, even though the agent calculates rewards internally, the reference itself is externally provided, so it ultimately has an extrinsic nature. Therefore, these approaches do not fully resolve the fundamental question of "Where does value come from, and can the agent establish it by itself?"
Reward-auxiliary Intrinsic Motivation
To allow the agent to establish its own value function independently, neither external rewards nor internal setpoints can be predetermined. But without these metrics, it is difficult to track learning progress or judge the utility of behaviors, which is why many studies on intrinsic motivation still tend to analyze it primarily as a means to facilitate exploration for extrinsic tasks (aubret_information-theoretic_2023). That is, the process by which agents generate and maintain goals themselves has been relatively less illuminated.
Heterostatic Intrinsic Motivation
Some works do investigate the behaviors arising purely from intrinsic motivation and qualitatively evaluate them in scenarios with no externally defined reward (eysenbach_diayn_2018; pathak_icm_2017). While these highlight the agent’s acquisition of diverse skills or behaviors, they typically involve a heterostatic approach, continuously seeking novel states. Consequently, once uncertainty is sufficiently reduced, further exploration tends to diminish, and fewer new behaviors emerge.
Active inference offers a unified theoretical framework that can integrate both extrinsic and intrinsic motivations under the single quantitative measure of free energy (parr_active_2022). Furthermore, (biehl_expanding_2018) showed that many existing methods for intrinsic motivation can be cast within active inference as processes for improving the posterior, a perspective reminiscent of (schmidhuber_formal_2010), who defined intrinsic motivation in terms of seeking “better world models.” However, such posterior-improving motives are likewise heterostatic: they promote exploration until uncertainty is resolved, after which they elicit fewer notable actions.
Homeostatic Intrinsic Motivation
Recently, studies have begun to emerge that address homeostatic intrinsic motivation, in which agents autonomously define their own setpoints but aim to preserve already learned states or past experiences, rather than continuously seeking novelty. These studies have made important contributions by demonstrating that goal-directed behaviors can emerge without external rewards. For example, (marcel_learning_2022) proposed a process in which latent representations of self-touch observations are used as setpoints to induce reaching behavior, and (takemura_neural_2018; kim_amam_2023) demonstrated that forward and inverse models learned from motor babbling can generate reaching toward target positions without explicit rewards. Most similar to our approach, (sajid_exploration_2021) introduced the concept of "learned preferences" within the free energy framework and semantically analyzed how a drive to reduce mismatches between observation and preference can lead to new behaviors. While these studies have demonstrated the potential of homeostatic intrinsic motivation, they operate within independent frameworks focused on a single intrinsic motivation mechanism, and integration with extrinsic motivations remains a future challenge.
Summary
In summary, our approach is differentiated from existing research in the following ways:
  1. 1.
    Intrinsic: The agent autonomously establishes criteria based on its own experience rather than adhering to externally fixed criteria.
  2. 2.
    Independent behavior generation: It independently generates specific goal-directed behaviors such as reaching, rather than serving an auxiliary role to facilitate exploration.
  3. 3.
    Homeostatic: Even in situations without information gain, behaviors persist without stopping as long as there is a mismatch with familiar states.
  4. 4.
    FEP with active inference: Defined within the active inference framework, it can be theoretically integrated with existing extrinsic motivations and heterostatic intrinsic motivations (information gain).
5.2 Origin of Early Intentional Behaviors
Infants initially depend heavily on reflex mechanisms for action. Over time, however, they gradually develop what (mele_intentional_1994) describes as “intentional behavior,” comprising (i) a motivational system that judges how to handle incoming stimuli and (ii) the capability to execute genuinely goal-directed actions. According to (zaadnoordijk_origins_2020), stimulus-driven intentions arise first, and then, with accumulating experience, more endogenous intentions emerge. For instance, newborns rely on reflexes such as rooting to acquire feeding experience; later, merely seeing a bottle prompts them to reach out and drink (stimulus-driven intention). As they become aware of internal states like hunger, they begin actively searching for the bottle themselves (endogenous intention).
Over time, to recognize themselves as independent entities, infants must undergo abundant experiences featuring multimodal redundancy and temporal–spatial contingencies (rochat_self-perception_1998). Sequentially developing intentional actions may spontaneously provide opportunities for such experiences. (thelen_transition_1993) similarly noted that infants initially lack both the ability and the intention to perform reaching, and that the formation of stable intentions, coupled with active exploration driven by those intentions, facilitates the learning of reaching.
However, prior studies have not discussed how intentions, including stimulus-driven intentions, and intentional behaviors develop through what structure. Even (priorelli_flexible_2023), which implemented switchable intentions within the active inference framework, used intentions that were discretely predefined rather than self-developing. Meanwhile, neurophysiological evidence suggests that the anterior cingulate cortex primarily deals with reward-related errors in individual decisions, whereas the posterior cingulate cortex processes discrepancy signals related to autobiographical information such as one’s own body or past experiences, and triggers new behaviors (pearson_posterior_2011; brewer_what_2013).
Our proposed self-prior aligns with this notion of “self-relevant predictions,” indicating that when new input fails to satisfy these predictions, the system initiates action. In other words, even without any externally designed goal, an intrinsic drive to preserve a self-prior learned from past experiences can manifest behaviors akin to an infant persistently touching a newly discovered sticker on their arm. This mirrors Piaget’s concept of primary circular reactions, where repeated attempts can refine intentions further, offering a theoretical foundation for computational studies on the origins of infant intentional behavior.
In our experiments, we fixed the self-prior’s learning rate. However, an excessively high rate could lead to overreactions (compulsive-like responses) to minor changes, whereas an overly low rate might cause indifference (apathy) toward new stimuli. Future work could investigate how more flexible learning rates shape different behavioral styles via self-relevant prediction-error processing.
5.3 Limitations and Future Directions
Our study employs a constructivist approach, focusing on demonstrating how the self-prior mechanism operates under minimal conditions. The infant sticker-reaching task serves as one example to clearly illustrate this mechanism, showing what behaviors emerge under the specific condition of tactile and proprioceptive body reaching. Since the self-prior is modality-independent in principle, the same mechanism can be applied to various sensory modalities. For example, using visual modality would be expected to produce reaching behavior toward rarely seen external objects outside the agent’s body. Indeed, several phenomena observed in actual infant development may also be explained by the self-prior mechanism. For instance, according to (eskandari_effect_2020), preterm infants nested in boundaries showed a significant reduction in unstable extension movements and maintained stable flexed postures, which can be reinterpreted as the infants preferring uterine-like environments and maintaining stable behavioral states to resolve mismatches with a self-prior learned from the flexed posture and tactile boundary sensations experienced during fetal stages.
Furthermore, behaviors that emerge in more complex cognitive contexts beyond simple sensorimotor-level adjustments, such as attention or social interactions, may also be explained by the self-prior. (warneken_altruistic_2006) presented tasks in which infants help adults who are having difficulty achieving goals, including results where infants opened a door for an adult who was having difficulty putting a magazine into a closet. Reinterpreting this in terms of the self-prior approach, it is possible that the infant opened the door to resolve a mismatch with a self-prior learned from visual information that adults normally open the closet door and put magazines in. Thus, the same principle of learning the probability density over experienced observations can explain behaviors across diverse contexts. Our approach has only been validated in discrete environments and low-degree-of-freedom (3-DoF) continuous environments, but real infants or robots exhibit far more complex and varied sensorimotor interactions. Therefore, it is necessary to extend this study to high-dimensional state and action environments. In particular, by utilizing simulations of early human sensorimotor experiences including fetal stages (kim_simulating_2022), it would be possible to explore how rich sensorimotor inputs provided in utero shape the self-prior and goal-directed behaviors, thereby revealing what behaviors can develop under physical constraints (kuniyoshi_fusing_2019).
In such complex environments, achieving generalizations such as reusing priors acquired across diverse contexts would require integrating all available modalities, learning various scenes over extended periods, and employing high-dimensional self-prior models capable of probabilistically modeling all of these. Such models could potentially be achieved through hierarchical active inference (friston_graphical_2017), or deep generative models like Transformers for handling long time series (vaswani_transformer_2017; chen_transdreamer_2022). Since the self-prior learns experienced sensations in a density model without prior knowledge, such extensions are possible in principle through scaling up generative models.
For example, generalizations such as transferring reaching behavior learned through touch to visual stimuli, generalizing responses to a specific shape of sticker to other shapes, or extending behaviors toward objects attached to one’s own body to external objects are all expected to be possible through large-scale generative models capable of learning statistical regularities across diverse sensory experiences. Such generalization can be understood not simply as repetition of the same action, but as a process of recognizing and utilizing structural similarities among self-priors formed in different contexts. Additionally, we showed behaviors driven solely by familiarity based intrinsic motivation by excluding externally imposed reward criteria, but in reality, infants or robots must balance multiple motivations such as energy homeostasis and novel information exploration. Therefore, designing experiments that systematically assess how these motivations complement or conflict and organize behavior in large-scale tasks will be an important task for future research.
Meanwhile, our study demonstrated that, by separating an intrinsic homeostatic term using a learnable self-prior from the extrinsic term using a fixed preferred prior that was commonly provided in prior studies using expected free energy, an agent can not only generate actions aimed at achieving predefined goals but also spontaneously generate goal-directed behaviors. However, a current limitation is that while the use of the self-prior is integrated into policy generation via expected free energy, the learning of the self-prior is performed separately from the learning of variational free energy.
One potential clue to resolving this separation may also lie in a hierarchical architecture. Specifically, a higher-level hidden state that reflects long-term experiences may be gradually learned, and the likelihood distribution inferred from it could serve as the self-prior for the lower level. This differs from conventional hierarchical inference, where the higher level simply learns priors and determines actions on a slower timescale to influence the lower level (e.g., $P(s_{t}|s_{t-1},a_{t-1})=\text{Cat}(\mathbf{B}_{a_{t-1}})$). Instead, it continuously receives information from the lower level to learn a prior, serving only to form a self-prior over a long duration (e.g., $P(s)=\text{Cat}(\mathbf{B})$). This hierarchical approach is also expected to be closely connected to building preferences over internal states. While the current implementation learns the self-prior in observation space, through a hierarchical approach, there is a possibility that higher-level internal latent representations may be formed through experience.
Another possibility is that the learning of the self-prior is handled not by the cortex but by a separate memory system such as the hippocampus, and thus may require structural separation from the inference system proposed under the free energy principle. Concrete validation of these open hypotheses is left for future work.
Finally, in the discrete environment, we could observe that as time passed with a sticker attached, the self-prior gradually changed and, accordingly, behavior also changed. By contrast, in the continuous environment, we conducted the experiment by fixing the probability of stickers being included in the observations used for self-prior learning at 5%. This was because the self-prior model stored past data in a replay buffer and learned without considering temporal order, making it difficult to investigate gradual changes in the self-prior according to changes in observation frequency. To reproduce the process by which stimulus-driven intention gradually strengthens into a more endogenous form of intention in real infants or robots, a self-prior learning mechanism that considers temporal order is necessary. This could potentially be implemented through the hierarchical architecture mentioned earlier, where higher levels integrate long-term experiences.
6 Conclusion
Building on the free energy principle and the active inference framework, we have demonstrated that an agent’s intrinsic drive to learn and maintain a "self-prior", which is akin to a body schema, can induce goal-directed behaviors even in the absence of externally specified rewards. In particular, whereas the conventional active inference literature has emphasized heterostatic intrinsic motivation (i.e., seeking new information), our work highlights a homeostatic form of intrinsic motivation, one that actively strives to preserve familiar sensory experiences.
Using a simulated infant touching and examining a sticker on its arm as an illustrative example, we showed how an internally derived drive to resolve mismatches between the self model and incoming data can manifest in behaviors such as reaching and sticker removal, all without any explicit reward criteria. This offers a computational interpretation of stimulus-driven intentional actions in developmental psychology and suggests potential relevance to the neurobiological literature on posterior cingulate cortex, which processes errors linked to autobiographical information.
Future work should extend our approach to complex physical environments such as full-scale infant simulations or high-DoF robots to investigate the long-term formation and refinement of intentions and behaviors. Additionally, we must experimentally validate how multiple motivations interact within a single free energy framework, incorporating not only familiarity based motivation but also extrinsic and information-seeking motivations. Such research will play an important role in analyzing the spontaneous developmental trajectory of intentional behavior.
Acknowledgments and Disclosure of Funding
This work was supported by JST, PRESTO Grant Number JPMJPR23S4, Japan.
Appendix A Appendix
A.1 Detailed Description of the Model for the Discrete Environment
A.1.1 Variational Free Energy Derivation
In the discrete model, observation $o_{t}$ and hidden state $s_{t}$ are one-hot column vectors with $N_{o}$ and $N_{s}$ rows, respectively. Categorical distribution parameters are implemented as matrices: $\mathbf{A}$ is an $N_{o}\times N_{s}$ matrix, and each $\mathbf{B}_{a}$ is an $N_{s}\times N_{s}$ square matrix. For instance, if actions are {LEFT, STOP, RIGHT}, separate matrices $\mathbf{B}_{\text{LEFT}},\mathbf{B}_{\text{STOP}},\mathbf{B}_{\text{RIGHT}}$ exist.
The variational free energy $\mathcal{F}$ expands as:
$\mathcal{F}$ $=\mathbb{E}_{Q(s_{t})}[\log Q(s_{t})-\log P(o_{t},s_{t})]$ (14)
$=\mathbb{E}_{Q(s_{t})}[\log{Q(s_{t})}-\log{P(s_{t})}]-\mathbb{E}_{Q(s_{t})}[\log P(o_{t}\mid s_{t})]$
$=\phi_{t}\cdot(\log{\phi_{t}}-\log{(\mathbf{B}_{a_{t-1}}\phi_{t-1})})-\phi_{t}\cdot(\log{(\mathbf{A}\cdot o_{t})})$
The parameter $\phi$ that minimizes free energy is derived by differentiation:
$\frac{\partial\mathcal{F}}{\partial\phi_{t}}$ $=\log{\phi_{t}}-\log{(\mathbf{B}_{a_{t-1}}\phi_{t-1})}+1-\log{(\mathbf{A}\cdot o_{t})}=0$ (15)
$\therefore\log{\phi_{t}}$ $\approx\log{(\mathbf{A}\cdot o_{t})}+\log{(\mathbf{B}_{a_{t-1}}\phi_{t-1})}$
$\phi_{t}$ $\approx\sigma(\log{(\mathbf{A}\cdot o_{t})}+\log{(\mathbf{B}_{a_{t-1}}\phi_{t-1})})$
This corresponds to Eq. (7) in the main text.
A.1.2 Expected Free Energy Derivation
The expected free energy can be derived through step-by-step approximations into a computationally tractable form:
$\mathcal{G}$ $=\mathbb{E}_{Q(s_{t+1},o_{t+1}\mid a_{t})}[\log Q(s_{t+1}\mid a_{t})-\log\tilde{P}(o_{t+1},s_{t+1}\mid a_{t})]$ (16)
$=\mathbb{E}_{Q(s_{t+1},o_{t+1}\mid a_{t})}[\log Q(s_{t+1}\mid a_{t})-\log P(s_{t+1}\mid o_{t+1},a_{t})-\log\tilde{P}(o_{t+1})]$
$\approx\mathbb{E}_{Q(s_{t+1},o_{t+1}\mid a_{t})}[\log Q(s_{t+1}\mid a_{t})-\log Q(s_{t+1}\mid o_{t+1},a_{t})-\log\tilde{P}(o_{t+1})]$
$=\mathbb{E}_{Q(s_{t+1},o_{t+1}\mid a_{t})}[\log Q(o_{t+1}\mid a_{t})-\log Q(o_{t+1}\mid s_{t+1},a_{t})-\log\tilde{P}(o_{t+1})]$
The approximation in the third line replaces the true posterior $P(s_{t+1}\mid o_{t+1},a_{t})$ with the approximate posterior $Q(s_{t+1}\mid o_{t+1},a_{t})$. The fourth line is derived by applying Bayes’ rule: $Q(s,o)=Q(o|s)Q(s)=Q(s|o)Q(o)$. Reformulating for computation:
$\mathcal{G}$ $\approx\mathbb{E}_{Q(s_{t+1}\mid a_{t})Q(o_{t+1}\mid s_{t+1},a_{t})}[-\log Q(o_{t+1}\mid s_{t+1},a_{t})]$ (17)
$\phantom{=======}+\mathbb{E}_{Q(o_{t+1}\mid a_{t})}[\log Q(o_{t+1}\mid a_{t})-\log\tilde{P}(o_{t+1})]$
$=\mathbb{E}_{Q(s_{t+1}\mid a_{t})P(o_{t+1}\mid s_{t+1})}[-\log P(o_{t+1}\mid s_{t+1})]$
$\phantom{=======}+\mathbb{E}_{Q(o_{t+1}\mid a_{t})}[\log Q(o_{t+1}\mid a_{t})-\log\tilde{P}(o_{t+1})]$
$=\mathbb{E}_{Q(s_{t+1}\mid a_{t})}[\mathcal{H}[P(o_{t+1}\mid s_{t+1})]]+D_{\mathrm{KL}}[Q(o_{t+1}\mid a_{t})\|\tilde{P}(o_{t+1})]$
$=(\mathbf{B}_{a_{t}}\phi_{t})\cdot\mathcal{H}[\mathbf{A}]+D_{\mathrm{KL}}[\mathbf{A}\mathbf{B}_{a_{t}}\phi_{t}\|\mathbf{C}]$
This corresponds to Eq. (9) in the main text. In our setup, since we assume that $\mathbf{A}$ and $\mathbf{B}$ are already known accurately, the action entropy term for information gain about parameters is omitted in the discrete environment experiment (see (parr_active_2022) for more details on the above equations).
A.2 Detailed Description of the Model for the Continuous Environment
A.2.1 Embedding Transformation Architecture for Continuous Environment Model
To represent both tactile and proprioceptive modalities as a unified observation $o_{t}$ in the model for the continuous environment, we convert each modality into an embedding of the same dimensionality. The tactile matrix ($80\times 30$) is transformed into an embedding using a Conv2D module, and the proprioceptive vector (3-dimensional) is transformed into an embedding using a MLP. We combine the two embeddings via element-wise addition to yield the final integrated observation embedding $o_{t}$. To decode from the embedding back to the original sensations, we use a ConvTranspose2D for tactile and a MLP for proprioception.
A.2.2 RSSM Computation
The deterministic state $h_{t}$ is computed as follows:
  1. 1.
    Encode previous stochastic state $s_{t-1}$ and action $a_{t-1}$ into a single vector via a linear layer
  2. 2.
    Feed the encoded vector together with previous deterministic state $h_{t-1}$ into a GRU cell (chung_empirical_2014)
  3. 3.
    The GRU cell output becomes the new deterministic state $h_{t}$
The stochastic prior $p_{\phi}(\hat{s}_{t}\mid h_{t})$ passes $h_{t}$ through a MLP to output mean and variance parameters of a multivariate Gaussian, from which the stochastic state is sampled.
The stochastic posterior $q_{\phi}(s_{t}\mid h_{t},o_{t})$ receives a concatenated vector of deterministic state $h_{t}$ and embedded observation $o_{t}$. The concatenated vector is passed through a MLP to output mean and variance parameters of a multivariate Gaussian, from which the stochastic state is sampled.
The likelihood $p_{\phi}(o_{t}\mid h_{t},s_{t})$ transforms the concatenated $h_{t}$ and $s_{t}$ into an embedded $o_{t}$ through a linear layer, which is then decoded into individual modality-specific observations.
All parameters $\phi$ constituting this RSSM structure are optimized via gradient descent to minimize the following variational free energy:
$\mathcal{F}$ $=\mathbb{E}_{q(s_{t})}[\log q(s_{t})-\log p(o_{t},s_{t})]$ (18)
$=\mathbb{E}_{q(s_{t})}[\log{q(s_{t})}-\log{p(s_{t})}]-\mathbb{E}_{q(s_{t})}[\log p(o_{t}\mid s_{t})]$
$=D_{\mathrm{KL}}[q(s_{t})\|p(s_{t})]-\mathbb{E}_{q(s_{t})}[\log p(o_{t}\mid s_{t})]$
$\therefore\operatorname*{\arg\min}_{\phi}\mathcal{F}$ $=\operatorname*{\arg\min}_{\phi}\big[D_{\mathrm{KL}}[q_{\phi}(s_{t}\mid h_{t},o_{t})\|p_{\phi}(s_{t}\mid h_{t})]-\mathbb{E}_{q_{\phi}(s_{t}\mid h_{t},o_{t})}[\log p_{\phi}(o_{t}\mid h_{t},s_{t})]\big]$
A.2.3 Policy and Value Network Training
Both the policy network $q_{\theta}(a_{t}\mid s_{t})$ and utility (value) network $g_{\psi}(s_{t})$ take the stochastic state $s_{t}$ as input, use a MLP to output multivariate Gaussian parameters, and sample from this distribution.
The training objectives are:
$\operatorname*{\arg\min}_{\theta}\mathcal{L}_{\text{policy}}$ $=\operatorname*{\arg\min}_{\theta}\sum\nolimits_{t}G^{\lambda}_{t}$ (19)
$\operatorname*{\arg\min}_{\psi}\mathcal{L}_{\text{utility}}$ $=\operatorname*{\arg\min}_{\psi}\sum\nolimits_{t}(g_{\psi}(s_{t})-G^{\lambda}_{t})$
[…]
Here, $G^{\lambda}_{t}$ is the GAE($\lambda$)-estimated expected utility approximation. $\gamma_{t}$ is the discount factor and $H$ is the simulation horizon.
A.3 Hyperparameters for Continuous Environment Experiments
Table 3 presents all hyperparameters used in the continuous environment experiments.
Table 3: Hyperparameters for continuous environment experiments
Parameter Value Description
Environment & Data Collection
Episode length ($L$) $1000$ Time steps per episode
Initial random episodes $100$ Random episodes before training
Exploration noise $0.3$ Noise added to policy during training
Training Schedule
Replay buffer size $450$ Max episodes stored in memory
Training epochs per episode $100$ Training iterations after each episode
Batch size ($B$) $50$ Number of trajectories sampled per epoch
Trajectory length ($L$) $50$ Time steps per sampled trajectory
Planning horizon ($H$) $15$ Future prediction horizon for policy learning
Gradient clipping $100.0$ Maximum gradient norm
Network optimizer AdamW Optimizer for training
Observation Embedding
CNN encoder channels $4,8,16,16$ Output channels for each encoder layer
CNN decoder channels $16,8,4,3$ Output channels for each decoder layer
CNN encoder kernel sizes $4,4,4,2$ Kernel size for each encoder layer
CNN decoder kernel sizes $2,4,4,4$ Kernel size for each decoder layer
CNN stride 22 Stride value for all layers
MLP hidden layer size $32$ Units in MLP hidden layer
MLP hidden layers 22 Number of hidden layers in MLP
Integrated embedding size $64$ Final embedding dimension
Activation layer ELU Activation layer
World Model (RSSM)
Deterministic state size ($\dim(h_{t})$) $200$ Dimension of GRU hidden state
Stochastic state size ($\dim(s_{t})$) $30$ Dimension of stochastic latent variable
Hidden layer size $200$ Units in hidden layers
Number of hidden layers 11 Number of hidden layers
Activation layer ELU Activation layer
Learning rate ($\alpha_{\phi}$) $10^{-3}$ Learning rate for world model parameters
Free nats $3.0$ Minimum threshold for KL divergence
KL scale ($\beta$) $2.0$ KL divergence weight ($\beta$-VAE)
Policy & Value Networks
Hidden layer size $200$ Units in hidden layers
Number of hidden layers 33 Number of hidden layers
Activation layer ELU Activation layer
Learning rate ($\alpha_{\theta},\alpha_{\psi}$) $10^{-3}$ Learning rate for parameters
Entropy temperature $10^{-4}$ Entropy regularization coefficient
Discount factor ($\gamma$) $0.99$ Discount rate for future rewards
GAE lambda ($\lambda$) $0.95$ Smoothing parameter for GAE estimation
Self-Prior (Neural Spline Flow)
Number of spline transforms 33 Number of transform layers in the flow
Hidden layer size $64$ Units in spline network hidden layers
Activation layer ReLU Activation layer
Learning rate ($\alpha_{\xi}$) $10^{-3}$ Learning rate for self-prior parameters